### v1

# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

Abstract
[…]
1 Introduction
Evaluating large language models against expensive oracle labels (human preferences, expert audits, downstream KPIs) is prohibitively costly at scale. The standard workaround is to use cheap LLM judges as proxies: collect judge scores on every sample, reserve oracle labels for a small validation slice, and hope the correlation is strong enough. This approach is ubiquitous in practice but statistically unsound.
Three systematic failures.
Uncalibrated “LLM-as-judge” evaluation exhibits three failure modes that compound in production:
  1. 1.
    Preference inversion. Judge scores $S$ are treated as if they were rewards $Y$ on the same scale. They are not. Without calibration, higher $S$ can predict lower $Y$, meaning the optimization target is fundamentally misaligned with the evaluation target. The “You’re absolutely right!” phenomenon (sycophantic responses scoring high on early metrics but low on actual user value) is preference inversion in production.
  2. 2.
    Invalid confidence intervals. Naive CIs on uncalibrated judge scores achieve near-0% coverage: their 95% CIs capture the true value in 0/50 seeds. Calibration with Oracle-Uncertainty-Aware (OUA) inference dramatically improves coverage. (Figure 5: at 5% oracle fraction, calibration uncertainty contributes over half of total variance.)
  3. 3.
[…]
Related work.
Recent work recognizes that LLM judges require calibration (Lee2025LLMJudgeReporting), but existing methods address only the simplest setting: single-model accuracy with binary outcomes. Guerdan2025DoublyRobustLLM use doubly robust methods to address external validity when LLM judges are prompted with demographic personas, a problem complementary to our focus on judge calibration for policy ranking. The harder problem of ranking multiple policies with continuous rewards under distribution shift remains inadequately addressed: no existing method provides calibration, overlap diagnostics, and valid uncertainty quantification in a unified framework. Without such a solution, practitioners either pay for full oracle labeling or accept invalid comparisons. CJE fills this gap.
Causal Judge Evaluation (CJE).
We introduce CJE, a framework that addresses all three failures. CJE casts offline evaluation as calibrated surrogate estimation with three core components:
  • AutoCal-R (reward calibration): Mean-preserving isotonic regression from judge score $S$ to oracle labels $Y$, with an automatic two-stage fallback for covariate-dependent bias (e.g., response length).
  • SIMCal-W (Surrogate-Indexed Monotone Calibration for Importance Weights): Variance-optimal stacking of $S$-monotone, unit-mean weight candidates, stabilizing weights for improved ESS.
  • OUA inference: Delete-one-fold jackknife that propagates calibration uncertainty into confidence intervals, dramatically improving coverage.
These components instantiate a single design rule, Design-by-Projection (DbP), that encodes justified knowledge as model restrictions. Building on established semiparametric theory, we show that when justified knowledge defines a restricted statistical model, the efficiency bound in the restricted model is at most the bound in the baseline model (Theorem 2); with cross-fitting, our estimators attain the surrogate information bound.
Beyond variance control, CJE includes a mean transport test: for each target policy $\pi^{\prime}$ or time period, we test whether the mean residual $\mathbb{E}[Y-f(S,X)]$ is zero. This is a necessary condition for reusing a single calibration function across policies or over time; if it fails, surrogate-only evaluation is systematically biased. In our Arena benchmark, the base-trained calibration transports cleanly to clone, premium, and prompt-engineered policies, but fails for an adversarial “unhelpful” policy, where mean residuals reveal a 0.31-0.31 level shift.
The Arena benchmark.
We validate CJE on a large-scale benchmark: 4,961 Chatbot Arena prompts, five LLM policies (including an adversarial “unhelpful” policy), 13 estimators, and GPT-5 as oracle. Key findings:
  • Calibration works. The Direct Method with two-stage calibration achieves 94% pairwise ranking accuracy on average (vs. 38% for SNIPS). Naive CIs on uncalibrated scores achieve 0% coverage; calibration with OUA improves coverage to 85–87% (Direct) and 95–96% (stacked-DR).
  • OPE fails unexpectedly. We expected weight stabilization to enable OPE; it did not. Despite SIMCal-W boosting ESS from $1\%$ to $80\%$, calibrated IPS remains near-random (47% pairwise). CLE explains why: high ESS is necessary but not sufficient when the logger rarely visits target-typical regions.
  • DR does not dominate. We expected DR to outperform Direct by combining logged data with fresh draws. Instead, Direct slightly outperforms DR (94.3% vs. 94.1%): under low overlap, DR’s IPS component adds noise rather than information.
Contributions.
  1. 1.
    Framework: CJE unifies calibration, weight stabilization, and uncertainty quantification for surrogate-based LLM evaluation.
  2. 2.
    Methods: AutoCal-R (mean-preserving calibration), SIMCal-W (weight stabilization for improved ESS), and OUA inference (calibration-aware CIs).
  3. 3.
    Theory: The CLE diagnostic provides a novel explanation for why high ESS is insufficient for IPS-style estimation, formalizing a precision floor under limited overlap. The Design-by-Projection framework applies established semiparametric principles to justify CJE’s methods.
  4. 4.
[…]
  5. 5.
    Operations: A closed-form Square Root Law for optimal budget allocation (Appendix F), enabling practitioners to minimize variance subject to cost constraints by balancing evaluation and calibration uncertainty.
Readers primarily interested in deployment can focus on Section 3 (pipeline overview), Section 5.3 (six practical rules), and Table 5 (estimator selection guide).
2 Background and Setup
Setup & notation.
We observe i.i.d. logs $(X_{i},A_{i},S_{i})$ under a fixed logger $\pi_{0}(\cdot\mid X)$; $S=s(X,A)$ is a scalar judge score on every row, and a small i.i.d. oracle slice provides labels $Y$. For a candidate policy $\pi^{\prime}$, the sequence-level importance ratio is
$W_{\pi^{\prime},i}\;=\;\frac{\pi^{\prime}(A_{i}\mid X_{i})}{\pi_{0}(A_{i}\mid X_{i})}\;=\;\exp\!\big\{\log p_{\pi^{\prime}}(A_{i}\mid X_{i})-\log p_{\pi_{0}}(A_{i}\mid X_{i})\big\},$
computed via teacher forcing (TF). The target is the counterfactual value $V(\pi^{\prime})=\mathbb{E}[Y(\pi^{\prime})]$. We use the sample-mean-one normalization (SNIPS) when helpful.
The LLM evaluation regime.
Standard off-policy evaluation (OPE) assumes the target policy cannot be executed. LLM evaluation operates in a different regime: generating text is cheap, but grading it with oracle labels (human experts, downstream KPIs) is expensive. This creates three evaluation modes: (i) logs-only (IPS/SNIPS), when we cannot re-run the model; (ii) fresh draws (Direct), when we generate new responses but score them with a calibrated surrogate; and (iii) hybrid (DR), combining both. All three share the core challenge: learning $S\to Y$ calibration from a small oracle slice collected under $\pi_{0}$, then applying it to evaluate $\pi^{\prime}$. The calibration is always off-policy; the actions may or may not be.
OPE basics.
IPS/SNIPS estimate $V(\pi^{\prime})$ by reweighting logged outcomes (HorvitzThompson1952; Hajek1964Rejective; LiChuLangfordSchapire2011; SwaminathanJoachims2015CRM). The direct method (DM) plugs in $g(x)=\sum_{a}\pi^{\prime}(a\mid x)\,\hat{m}(x,a)$. Doubly robust (DR) estimators combine IPS and DM (BangRobins2005) and, with sample–splitting and cross–fitting, admit $\sqrt{n}$ inference under the standard one–of–two $n^{-1/4}$ product–rate condition (Bickel1993; VaartWellner2000; Kosorok2008; JiangLi2016; Chernozhukov2018; vanDerLaanRose2011TL; KallusUehara2020DRL). Teacher forcing (TF) provides sequence–level propensities/ratios, so these forms apply to sequence policies in principle; however, TF reliability depends on API implementation details (see Limitations).
Variance, overlap, and stabilization.
IPS variance scales with $\mathbb{E}[W_{\pi^{\prime}}^{2}]$ and deteriorates under limited overlap (Crump2009LimitedOverlap). We monitor stability with the effective sample size (ESS),
$\mathrm{ESS}(W)=\frac{\big(\sum_{i}W_{i}\big)^{2}}{\sum_{i}W_{i}^{2}},\qquad\frac{\mathrm{ESS}(W)}{n}=\frac{1}{1+\mathrm{CV}^{2}(W)}\ \ \text{when }\bar{W}=1\ \text{(global mean–one/SNIPS)}.$
We also track tail behavior via diagnostics (Hill1975TailIndex; Liu2001MC; Owen2013MCTME). Common stabilizers include truncation/clipping (Ionides2008TruncatedIS), overlap weighting (LiMorganZaslavsky2018), balancing objectives (Kallus2018Balanced), and covariate–shift reweighting (Shimodaira2000CovShift; SugiyamaKrauledatMueller2007IWCV).
The Coverage-Limited Efficiency (CLE) problem.
High ESS indicates that weights are not dominated by a few extreme observations, but it does not guarantee that the logger has meaningful coverage in regions where the target policy concentrates. Let $\mathcal{T}\subset\mathcal{X}\times\mathcal{A}$ be a target-relevant region with $\alpha=P_{\pi^{\prime}}(\mathcal{T})$ (target mass) and $\beta=P_{\pi_{0}}(\mathcal{T})$ (logger mass). For any IPS-style estimator based on importance-weighted outcomes,
$\mathrm{SE}(\hat{\Psi}_{\mathrm{IPS}})\;\geq\;\frac{\sigma_{\mathcal{T}}\,\alpha}{\sqrt{\beta\,n}}\sqrt{1+\chi^{2}\!\big(\pi^{\prime}_{\mathcal{T}}\,\|\,\pi_{0,\mathcal{T}}\big)},$ (1)
where $\sigma_{\mathcal{T}}^{2}:=\mathrm{ess\,inf}_{(x,a)\in\mathcal{T}}\mathrm{Var}(Y\mid X{=}x,A{=}a)$ is the minimal outcome noise in $\mathcal{T}$, and $\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})$ measures shape mismatch inside $\mathcal{T}$ (proof in Appendix C.12). The diagnostic has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions; (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, inflating the floor even with good coverage; and (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
[…]
Diagnostics for CLE.
The CLE bound decomposes into two factors, each with a corresponding diagnostic:
  • Coverage penalty ($\alpha/\sqrt{\beta}$): Measured by TTC (Target-Typicality Coverage) $\hat{\beta}$, the estimated logger mass on $\mathcal{T}$. We define the target-typical region $\mathcal{T}$ via a surprisal threshold: let $\ell_{\pi^{\prime}}(x,a)=-\log p_{\pi^{\prime}}(a\mid x)$ be the total negative log-likelihood under the target policy (from teacher forcing), and let $\bar{\ell}_{\pi^{\prime}}(x,a)=\ell_{\pi^{\prime}}(x,a)/|a|$ be the mean per-token surprisal. Set $\mathcal{T}=\{(x,a):\bar{\ell}_{\pi^{\prime}}(x,a)\leq q_{0.9}\}$, where $q_{0.9}$ is the 90th percentile of $\bar{\ell}_{\pi^{\prime}}$ under $\pi^{\prime}$. Then $\alpha=P_{\pi^{\prime}}(\mathcal{T})=0.9$ by construction, and TTC $\hat{\beta}=\hat{P}_{\pi_{0}}(\mathcal{T})$ is the empirical fraction of logged samples falling in $\mathcal{T}$. If TTC<0.7{}<0.7, logs-only IPS will fail, meaning the CLE lower bound implies standard errors too large to distinguish realistic policy differences, even under ideal calibration and weight stabilization; prefer Direct or DR methods. (Threshold calibrated for MDE ${\approx}0.02$ at $n{\approx}5\text{k}$; see App. D.7.) Applicability: TTC requires sampling from or computing surprisal under $\pi^{\prime}$ (as we do via teacher forcing). When $\pi^{\prime}$-access is unavailable, fall back to conservative rules (e.g., always prefer Direct over IPS) or use proxy distributions to approximate TTC.
  • Shape mismatch ($\sqrt{1+\chi^{2}}$): Measured by Bhattacharyya affinity $A_{B}=\int\sqrt{p_{S|\pi^{\prime}}(s)\,p_{S|\pi_{0}}(s)}\,ds$ in the surrogate space. Low affinity ($A_{B}<0.5$) indicates severe shape mismatch; the gate threshold ($A_{B}\geq 0.85$; Table 6) ensures adequate overlap for reliable IPS.
Both diagnostics are complementary: TTC checks if the logger visits target-typical regions (action space); Bhattacharyya affinity checks overall alignment of judge score distributions across policies (surrogate space).
Calibration for OPE.
Calibration enforces identities under $\pi_{0}$: outcome calibration de-biases $g(X)$; ratio calibration enforces $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}]=1$ and $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}h]=\mathbb{E}_{\pi^{\prime}}[h]$ for a test class $h$; orthogonal moments enable honest inference (Chernozhukov2018). Recent work gives projection-based IPS/DR with stability guarantees (vanDerLaanLinCaroneLuedtke2025StabIPW; vanDerLaanLuedtkeCarone2024DRCalibration). For DR, IF orthogonality renders small calibration error second order (Bickel1993; Chernozhukov2018; vanDerLaanRose2011TL).
Shape constraints (isotonic).
Isotonic regression is the Euclidean projection onto the cone of monotone functions (PAVA) (Ayer1955PAVA; Barlow1972); it avoids extrapolation and weakly reduces dispersion by majorization (HardyLittlewoodPolya1952; MarshallOlkinArnold2011). CJE exploits this via mean-preserving isotonic projections for both reward calibration and weight stabilization (Section 3).
Judges as surrogates.
Automatic judges (LLM-as-judge or preference models) provide scalable scoring (Ouyang2022Instruct; BaiEtAl2022; Zheng2023LLMasJudge; Kim2024PrometheusJudge; KocmiFedermann2023) but are correlational and may drift (DietzEtAl2025LLMJudges). Viewing $S$ as a surrogate connects to surrogate endpoints and mediation (Prentice1989Surrogate; RobinsGreenland1992; FrangakisRubin2002PS; Pearl2012Mediation; VanderWeele2015Book). The surrogate paradox, wherein a policy improves $S$ and $S$ correlates with $Y$ yet the policy harms $Y$, can arise even with high correlation (VanderWeele2013Surrogates); our transport test (3) provides a necessary check. Under mean sufficiency ($\mathbb{E}[Y\mid X,A,S]=\mu(S)$), calibrating $R=f(S)$ preserves $V(\pi^{\prime})=\mathbb{E}[f(S^{\pi^{\prime}})]$ and supplies a one-dimensional index that stabilizes weights.
OUA uncertainty & IF stacking.
Treating learned $R=\hat{f}(S)$ as fixed understates uncertainty when the oracle slice is small. CJE addresses this via Oracle-Uncertainty-Aware (OUA) inference, which propagates calibration uncertainty into confidence intervals using a delete-one-fold jackknife (EfronStein1981; Bickel1993). For ensemble estimation, CJE stacks multiple estimators by minimizing IF covariance over the simplex (Wolpert1992StackedGeneralization; vanderLaan2007SuperLearner). Details in Section 3.
3 Methods
The CJE pipeline has three components that directly address the three failures identified above: reward calibration to prevent preference inversion, weight stabilization to control IPS variance, and uncertainty-aware inference that dramatically improves CI coverage. Figure 1 provides an overview.
Figure 1: CJE pipeline overview. A small oracle slice (5–25%) provides expensive oracle labels to train a calibration model ($S\to Y$). The learned mapping is then applied to bulk evaluation data where oracle labels are unavailable, enabling policy evaluation at a fraction of the cost. (In experiments, the oracle is GPT-5; in production this would typically be human raters or downstream KPIs.)
CJE follows one design rule, Design-by-Projection (DbP), applied to each object in the pipeline: (i) calibrate the reward (projection onto a monotone cone), (ii) stabilize ratios (projection onto a unit-mean, $S$-monotone cone), (iii) compute an orthogonalized estimator (projection onto a nuisance-orthogonal subspace), and (iv) optionally hedge variance by stacking (projection onto a simplex). All learners are cross-fitted; by the Efficiency via Model Restriction theorem, these projections encode justified knowledge that lowers the efficiency bound (see Section 4).
3.1 Reward calibration (AutoCal-R: isotonic in $S$ with an automatic two-stage fallback)
On the oracle slice $\{(S_{i},Y_{i})\}$, fit a mean-preserving calibrator $R=f(Z(S))$ with $K$-fold cross-fitting:
  • Monotone mode (default). Isotonic regression on $S$: $\hat{f}_{\uparrow}\in\arg\min_{f\in\mathcal{M}_{\uparrow}}\sum_{i\in O}\big(Y_{i}-f(S_{i})\big)^{2}$. PAVA preserves the slice mean exactly.
  • Two-stage mode (automatic fallback). Fit a smooth index $Z(S,X)=g(S,X)$ (splines+ridge), map to mid-ranks $U=\mathrm{ECDF}\{Z(S,X)\}$, then fit isotonic $\hat{h}_{\uparrow}(U)$. Predictions are $R=\hat{h}_{\uparrow}(\mathrm{ECDF}\{g(S,X)\})$.
Remark (covariate flexibility). The terminal isotonic step preserves mean-honesty regardless of the first-stage model, so $g$ can be any cross-fitted learner (splines, random forests, gradient boosting) and can incorporate covariates $X$ beyond $S$. In our experiments, including response length as a first-stage covariate significantly improves ranking accuracy: LLM judges often favor longer responses independent of quality, and conditioning on length removes this confounder from the $S\to Y$ mapping. Select the mode by OOF RMSE with a one-standard-error (1-SE) preference for monotone. Let $R^{\mathrm{OOF}}$ denote OOF predictions used along the IF path; the point estimate may use the pooled fit. The terminal isotonic step makes AutoCal-R mean-honest in either mode. We treat $\hat{f}$ as learned and propagate its uncertainty via OUA inference (Section 3.5).
Figure 2: AutoCal-R: monotone vs. two-stage calibration. Left: standard isotonic regression enforces monotonicity but cannot capture non-monotonic patterns in $\mathbb{E}[Y\mid S]$. Right: two-stage calibration (spline index $\to$ isotonic) can fit flexible patterns while preserving the mean. AutoCal-R automatically selects the mode via cross-validation.
3.2 Weight calibration (SIMCal-W: unit-mean ratios via OOF stacking of $S$-monotone candidates)
Let $W_{\pi^{\prime}}^{\mathrm{m1}}$ be the sample-mean-one baseline (SNIPS). For each fold $k$:
  1. 1.
    Monotone projections (train on $I_{\neg k}$). Fit increasing and decreasing isotonic maps on $S$ (the latter by running PAVA on S-S), rescale each to mean one on $I_{\neg k}$, and predict OOF candidates on $I_{k}$: $W_{\uparrow}^{\mathrm{OOF}}$, $W_{\downarrow}^{\mathrm{OOF}}$; optionally include an identity candidate $W_{\mathrm{base}}^{\mathrm{OOF}}\equiv 1$.
  2. 2.
    OOF stacking (variance-aware). Define residuals $\Delta_{i}$ used by the downstream estimator: $\Delta_{i}{=}R_{i}$ for IPS and $\Delta_{i}{=}R_{i}-\hat{m}(X_{i},A_{i})$ for DR (with $\hat{m}{=}\hat{q}$ below). Let $U_{c}=W_{c}^{\mathrm{OOF}}\Delta$ for $c\in\{\mathrm{base},\uparrow,\downarrow\}$ and compute $\hat{\Sigma}_{cd}=\mathrm{cov}(U_{c},U_{d})$ (tiny ridge if needed). Choose simplex weights
    $\hat{\beta}\in\arg\min_{\beta\in\Delta_{3}}\ \beta^{\top}\hat{\Sigma}\,\beta,\qquad W^{\mathrm{stack}}=\sum_{c}\hat{\beta}_{c}\,W_{c}^{\mathrm{OOF}},$
    then renormalize $W^{\mathrm{stack}}$ to mean one.
  3. 3.
    Light variance guard (optional; $\rho{=}1$ by default). Cap variance at absolute threshold $\rho$:
    $\alpha=\min\!\Big\{1,\ \sqrt{\frac{\rho}{\operatorname{Var}(W^{\mathrm{stack}})}}\Big\},\quad W^{\mathrm{blend}}=1+\alpha\,(W^{\mathrm{stack}}-1),\quad\hat{W}_{\pi^{\prime}}=W^{\mathrm{blend}}/\bar{W}^{\mathrm{blend}}.$
    Each step preserves (or restores) the sample mean. Note: The theoretical guarantee (Lemma 5) uses a relative bound $\operatorname{Var}(\hat{W})\leq\rho\,\operatorname{Var}(W^{\mathrm{m1}})$; the implementation’s absolute cap is more conservative when $\operatorname{Var}(W^{\mathrm{m1}}){>}1$.
Remark (transport view). In the continuous case the ideal component is $m^{\star}(s)=p_{S\mid\pi^{\prime}}(s)/p_{S\mid\pi_{0}}(s)$; SIMCal-W stacks increasing and decreasing $S$-monotone candidates to minimize IF variance, yielding approximately monotone weights when one direction dominates.
[…]
3.3 Estimators: Calibrated IPS and DR-CPO
Let $\hat{q}(z)\approx\mathbb{E}[R\mid Z{=}z]$ be the cross-fitted calibrator (in experiments, $\hat{q}=\hat{f}^{(-k)}(Z)$; we do not fit a separate text-conditioned reward model). Define $\hat{g}_{\pi^{\prime}}(x)=\mathbb{E}_{\pi^{\prime}}[\hat{q}(Z)\mid X{=}x]$; all nuisances are cross-fitted and OOF predictions are used inside IFs.
Calibrated IPS.
$\widehat{V}_{\mathrm{IPS}}(\pi^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\hat{W}_{\pi^{\prime},i}\,R_{i},\qquad\phi^{\mathrm{IPS}}_{i}=\hat{W}_{\pi^{\prime},i}\,R^{\mathrm{OOF}}_{i}-\widehat{V}_{\mathrm{IPS}}.$
DR-CPO (Doubly Robust Calibrated Policy Outcomes).
$\widehat{V}_{\mathrm{DR}}(\pi^{\prime})=\frac{1}{n}\sum_{i=1}^{n}\Big\{\hat{g}_{\pi^{\prime}}(X_{i})+\hat{W}_{\pi^{\prime},i}\big(R_{i}-\hat{q}(Z_{i})\big)\Big\},\quad\phi^{\mathrm{DR}}_{i}=\hat{g}_{\pi^{\prime}}(X_{i})+\hat{W}_{\pi^{\prime},i}\big(R^{\mathrm{OOF}}_{i}-\hat{q}^{\mathrm{OOF}}_{i}\big)-\widehat{V}_{\mathrm{DR}}.$
3.4 IF-space stacking (variance-optimal convex ensembling)
For a small library $\mathcal{E}$ of regular estimators (e.g., DR/TMLE/MRDR variants, capped IPS), form the matrix of centered IF columns $\Phi=[\phi^{(e)}]_{e\in\mathcal{E}}$ (computed OOF on the same folds), estimate $\hat{\Sigma}=\tfrac{1}{n}\Phi^{\top}\Phi+\lambda I$, and solve the simplex QP
$\hat{\alpha}\in\arg\min_{\alpha\in\Delta}\ \alpha^{\top}\hat{\Sigma}\,\alpha,\qquad\widehat{V}_{\mathrm{stack}}=\sum_{e\in\mathcal{E}}\hat{\alpha}_{e}\,\widehat{V}^{(e)},\quad\phi^{\mathrm{stack}}=\sum_{e\in\mathcal{E}}\hat{\alpha}_{e}\,\phi^{(e)}.$
3.5 Oracle-Uncertainty-Aware (OUA) inference
Standard CIs treat the calibration function $\hat{f}$ as fixed, ignoring first-stage estimation error. When the oracle slice is small, this understates total variance and produces dramatically narrow intervals. In our experiments, naive CIs on uncalibrated scores achieve 0% coverage; calibration with OUA improves coverage to 85–96%.
OUA variance decomposition.
Total variance decomposes into two components:
[…]
where $\operatorname{Var}_{\text{eval}}$ is the standard evaluation variance (estimated via the IF) and $\operatorname{Var}_{\text{cal}}$ is the variance due to calibration uncertainty.
Delete-one-fold jackknife.
We estimate $\operatorname{Var}_{\text{cal}}$ via a delete-one-oracle-fold jackknife:
  1. 1.
    Partition the oracle slice into $K$ folds.
  2. 2.
    For each fold $k\in\{1,\ldots,K\}$, refit calibration omitting fold $k$: $\hat{f}^{(-k)}$.
  3. 3.
    Compute the leave-one-fold-out estimate: $\widehat{V}^{(-k)}=\widehat{V}(\hat{f}^{(-k)})$.
  4. 4.
    Estimate calibration variance:
    $\widehat{\operatorname{Var}}_{\text{cal}}=\frac{K-1}{K}\sum_{k=1}^{K}\big(\widehat{V}^{(-k)}-\bar{V}\big)^{2},\quad\bar{V}=\frac{1}{K}\sum_{k}\widehat{V}^{(-k)}.$
OUA confidence intervals.
$\widehat{\operatorname{Var}}_{\text{total}}=\widehat{\operatorname{Var}}_{\text{eval}}+\widehat{\operatorname{Var}}_{\text{cal}},\qquad\text{CI}_{95\%}=\widehat{V}\pm 1.96\sqrt{\widehat{\operatorname{Var}}_{\text{total}}}.$
The OUA share $\widehat{\operatorname{Var}}_{\text{cal}}/\widehat{\operatorname{Var}}_{\text{total}}$ ranges from 0% (100% oracle fraction, where no calibration is needed) to over 50% (5% oracle fraction), explaining why ignoring it produces catastrophically narrow CIs.
Computational cost.
OUA adds $K$ refits of AutoCal-R (isotonic regression) and $K$ re-evaluations of the estimator. Since both are $O(n\log n)$, the overhead is modest relative to the cost of teacher forcing and response generation.
4 Theory: EIF, Design-by-Projection, and Efficiency
This section explains why the projections in CJE are not heuristics but information-optimal. We show: (i) using a judge as a surrogate can only reduce variance, (ii) encoding justified restrictions like monotonicity and mean-one weights tightens efficiency bounds, and (iii) our estimators attain these bounds under standard conditions. Proofs and technical lemmas are deferred to the appendix.
Surrogate model and EIF.
Let $R^{\star}=\mathbb{E}[Y\mid S]$ and $m^{\star}(S)=\mathbb{E}[W_{\pi^{\prime}}\mid S]$. Under mean sufficiency $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$,
$V(\pi^{\prime})=\mathbb{E}\!\big[m^{\star}(S)\,R^{\star}(S)\big],\qquad\phi_{\mathrm{sur}}(O;\pi^{\prime})\;=\;g^{\star}_{\pi^{\prime},R}(X)\;+\;m^{\star}(S)\!\big(R^{\star}-q^{\star}_{R}(X,A)\big)\;-\;V(\pi^{\prime}),$
with $q^{\star}_{R}(x,a)=\mathbb{E}[R^{\star}\mid X{=}x,A{=}a]$ and $g^{\star}_{\pi^{\prime},R}(x)=\sum_{a}\pi^{\prime}(a\mid x)\,q^{\star}_{R}(x,a)$.
Theorem 1 (Surrogate EIF and variance reduction).
Let $\phi_{\mathrm{uncon}}$ be the canonical gradient in the nonparametric model that does not use $S$. Then $\phi_{\mathrm{sur}}$ is the canonical gradient in the surrogate model, and $\operatorname{Var}(\phi_{\mathrm{sur}})\leq\operatorname{Var}(\phi_{\mathrm{uncon}})$, with strict inequality unless $W_{\pi^{\prime}}$ is $\sigma(S)$-measurable and $R^{\star}$ is degenerate.
In plain terms: If the judge carries any information about the oracle, using it can only help variance and never hurt it. Coarser judges (garblings) are strictly worse. (This variance reduction from conditioning is a standard result; we state it to establish the foundation for subsequent results.)
Efficiency via Model Restriction.
Let $L^{2}_{0}$ be the mean-zero Hilbert space with inner product $\langle f,g\rangle=\mathbb{E}[fg]$. Let $\mathcal{M}$ be a baseline semiparametric model with tangent space $T(P)$ and canonical gradient (EIF) $\phi^{\star}$. When justified knowledge defines a restricted model $\mathcal{M}_{\mathcal{C}}\subset\mathcal{M}$ (e.g., mean sufficiency in $S$, monotonicity of conditional expectations), the restricted tangent space $T_{\mathcal{C}}(P)\subseteq T(P)$ yields a new canonical gradient $\phi^{\star}_{\mathcal{C}}$.
Theorem 2 (Efficiency via Model Restriction).
(i) Efficiency improvement. The canonical gradient in the restricted model satisfies $\|\phi^{\star}_{\mathcal{C}}\|_{2}^{2}\leq\|\phi^{\star}\|_{2}^{2}$, with equality iff $T_{\mathcal{C}}(P)=T(P)$. If $T_{\mathcal{C}_{1}}(P)\supseteq T_{\mathcal{C}_{2}}(P)$ (i.e., $\mathcal{C}_{2}$ imposes stronger restrictions), then $\|\phi^{\star}_{\mathcal{C}_{2}}\|_{2}^{2}\leq\|\phi^{\star}_{\mathcal{C}_{1}}\|_{2}^{2}$. (ii) Attainability. One-step/TMLE estimators with cross-fitting attain the efficiency bound $\operatorname{Var}(\phi^{\star}_{\mathcal{C}})$ in the restricted model $\mathcal{M}_{\mathcal{C}}$.
In plain terms: Every justified assumption we encode as a model restriction is “free variance reduction”: we shrink the tangent space without changing the estimand.
Remark (prior art). Theorem 2 formalizes a well-known principle in semiparametric efficiency theory (Newey1990Semiparametric; Bickel1993); our contribution is naming it and systematically applying it to surrogate-based LLM evaluation via the Design-by-Projection framework.
Corollary 1 (Blackwell–efficiency monotonicity).
If $S_{2}$ is a garbling of $S_{1}$ (i.e., $\sigma(S_{2})\subseteq\sigma(S_{1})$), then $\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{1})\big)\leq\operatorname{Var}\!\big(\phi_{\mathrm{sur}}(S_{2})\big)$, with strict inequality unless $W_{\pi^{\prime}}$ is already $\sigma(S_{2})$-measurable and $R^{\star}$ is degenerate.
Consequences for CJE
(i) Conditioning: taking $\mathcal{C}=\{f:\,f=\mathbb{E}[f\mid S]\}$ recovers Theorem 1. (ii) Mean-one monotone weights: restricting the weight component to $\{w:\mathbb{E}[w]=1,\;w\uparrow S\}$ motivates SIMCal-W; the exact $\mathrm{IsoMeanOne}_{S}$ projection weakly reduces dispersion in finite samples by majorization. (iii) Stacking: restricting to the convex hull of candidate IF columns gives the variance-optimal convex ensemble.
Proposition 1 (Cal-IPS: mean correctness and dispersion control).
Let $R=\hat{f}(Z(S))$ be AutoCal-R (monotone in $S$ or two-stage index; cross-fitted), and let $W^{\mathrm{m1}}_{\pi^{\prime}}\!\triangleq W_{\pi^{\prime}}/\bar{W}_{n}$ denote the sample-mean-one (SNIPS) ratios, where $\bar{W}_{n}=n^{-1}\sum_{i}W_{\pi^{\prime},i}$. Let $\hat{W}_{\pi^{\prime}}$ be SIMCal-W weights (OOF stack ++ variance guard $\rho\geq 1$). Then $\widehat{V}_{\mathrm{IPS}}=\tfrac{1}{n}\sum_{i}\hat{W}_{\pi^{\prime},i}R_{i}\to_{p}V(\pi^{\prime})$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection (see App. A.3), $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ and $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$; the reference implementation uses a simpler normalization that empirically yields similar ESS gains.
Theorem 3 (DR-CPO: $\sqrt{n}$ limits and efficiency).
Assume mean sufficiency, suitable tails/moments, and cross-fitted nuisances satisfying the one-of-two rate condition $\|\hat{q}-q^{\star}_{R}\|_{2}\cdot\|\hat{W}_{\pi^{\prime}}-m^{\star}\|_{2}=o_{p}(n^{-1/2})$ (e.g., either factor $o_{p}(n^{-1/4})$). Then
$\sqrt{n}\big(\widehat{V}_{\mathrm{DR}}-V(\pi^{\prime})\big)\ \leadsto\ \mathcal{N}\!\big(0,\operatorname{Var}(\phi_{\mathrm{sur}})\big),$
i.e., DR-CPO attains the surrogate efficiency bound. (Efficiency attainment under the product-rate condition is a standard property of DR estimators (BangRobins2005; Chernozhukov2018); our contribution is verifying that the CJE construction satisfies these conditions.)
Budgeted bound (variance cap).
For $\rho\geq 1$, define
$\mathcal{W}_{\rho}\;=\;\Big\{m:\ \mathbb{E}[m]=1,\;m\uparrow S,\;\mathbb{E}\big[(m-1)^{2}\big]\leq\rho\,\mathbb{E}\big[(m^{\star}-1)^{2}\big]\Big\},$
and let $m^{\star}_{\rho}$ be the $L^{2}(P_{S})$ projection of $m^{\star}$ onto $\mathcal{W}_{\rho}$. Define the budgeted gradient $\phi^{(\rho)}=g^{\star}_{\pi^{\prime},R}(X)+m^{\star}_{\rho}(S)\big(R^{\star}-q^{\star}_{R}(X,A)\big)-V(\pi^{\prime})$.
Theorem 4 (Budgeted information bound).
The optimal asymptotic variance under the cap $\rho$ equals $\operatorname{Var}(\phi^{(\rho)})$, which is nonincreasing in $\rho$ and satisfies $\lim_{\rho\to\infty}\operatorname{Var}(\phi^{(\rho)})=\operatorname{Var}(\phi_{\mathrm{sur}})$. If SIMCal-W converges to $m^{\star}_{\rho}$ (with the guard), then DR-CPO attains $\operatorname{Var}(\phi^{(\rho)})$.
In plain terms: This formalizes “don’t let weights explode.” Limiting weight variance ($\rho$) trades a bit of asymptotic efficiency for much better finite-sample behavior.
Theorem 5 (IF-space stacking).
Let $\{\widehat{V}^{(e)}\}_{e\in\mathcal{E}}$ be regular, asymptotically linear estimators with centered IFs $\{\phi^{(e)}\}$, and let $\hat{\alpha}\in\arg\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}$ the empirical IF covariance. If $\hat{\Sigma}\to\Sigma$ uniformly on $\Delta$, then the stacked estimator $\widehat{V}^{(\hat{\alpha})}$ is asymptotically linear with IF $\phi^{(\alpha^{\star})}$ and variance $\min_{\alpha\in\Delta}\alpha^{\top}\Sigma\,\alpha\leq\min_{e}\Sigma_{ee}$. An outer split leaves the limit unchanged. (This extends ensemble learning from prediction, where Super Learner (vanderLaan2007SuperLearner) combines algorithms to minimize cross-validated risk, to asymptotic efficiency: we combine estimators by minimizing the variance of their combined influence function.)
Corollary 2 (Carathéodory sparsity).
If $\Phi=[\phi^{(e)}]$ has empirical rank $r$, the variance-optimal convex combination uses at most r+1r{+}1 base estimators.
Proposition 2 (OUA jackknife variance estimation).
[…]
Policy-wise mean transport test.
Let $f(S,X)$ be a calibration function learned on the oracle slice under the logging policy $\pi_{0}$. For a target policy $\pi^{\prime}$, define the oracle and surrogate values $V_{\pi^{\prime}}^{\mathrm{oracle}}:=\mathbb{E}_{\pi^{\prime}}[Y]$ and $V_{\pi^{\prime}}^{\mathrm{sur}}:=\mathbb{E}_{\pi^{\prime}}[f(S,X)]$, and the residual $\varepsilon_{\pi^{\prime}}:=Y-f(S,X)$.
Proposition 3 (Mean transport equivalence).
$V_{\pi^{\prime}}^{\mathrm{oracle}}-V_{\pi^{\prime}}^{\mathrm{sur}}=\mathbb{E}_{\pi^{\prime}}[\varepsilon_{\pi^{\prime}}]$. Therefore, the null hypothesis $H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[\varepsilon_{\pi^{\prime}}]=0$ is equivalent to mean-unbiasedness of the surrogate for policy $\pi^{\prime}$: $H_{0,\pi^{\prime}}\iff V_{\pi^{\prime}}^{\mathrm{oracle}}=V_{\pi^{\prime}}^{\mathrm{sur}}$.
Remark. The mean residual decomposes into a transportability gap (failure of S2: $\mathbb{E}_{\pi^{\prime}}[\mu_{\pi^{\prime}}(S,X)-\mu_{\pi_{0}}(S,X)]$) and calibration estimation error ($\mathbb{E}_{\pi^{\prime}}[\mu_{\pi_{0}}(S,X)-f(S,X)]$). The latter vanishes as the oracle slice grows, so the test asymptotically isolates transportability failures. This test is strictly weaker than full transportability (S2): it guarantees mean-unbiasedness for policy-wise values but does not rule out conditional miscalibration within subgroups or at tails.
Data requirement. Applying this test requires a small oracle slice under each target policy $\pi^{\prime}$ (or per evaluation environment, e.g., time window or subgroup). We recommend collecting a modest audit sample (${\sim}50$200200 labels per policy) to validate that the base-trained calibrator can be safely reused; if a policy fails the test, either recalibrate with policy-specific data or fall back to oracle-only evaluation for that cell.
Discussion.
Theorem 2 formalizes Design-by-Projection: when justified knowledge defines a restricted statistical model (smaller tangent space), the efficiency bound in the restricted model is at most the bound in the baseline model. The key instantiations are: Theorem 1 (conditioning on $S$), Theorem 4 (budgeted monotone weights), and Theorem 5 (convex hull of candidate IFs). AutoCal-R, SIMCal-W, and DR/TMLE with cross-fitting attain the corresponding bounds in their respective restricted models. In finite samples, the exact $\mathrm{IsoMeanOne}_{S}$ projection additionally majorizes dispersion (Lemma 2), explaining the ESS gains delivered by SIMCal-W.
Scope and limitations.
(i) Mean sufficiency generalization. Theorems 1 and 3 assume mean sufficiency in $S$: $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$. AutoCal-R’s two-stage mode uses a learned index $Z(S,X)$; the theory generalizes by replacing $S$ with $Z(S,X)$ throughout, provided $Z$ is consistently estimated. (ii) Double robustness and weight misspecification. Theorem 3 assumes at least one nuisance converges at rate $n^{-1/4}$. When teacher-forced weights are unreliable (as in high-overlap-failure regimes), DR remains consistent via the outcome model alone; however, the efficiency bound $\operatorname{Var}(\phi_{\mathrm{sur}})$ is attained only when both nuisances are well-specified. In practice, when IPS fails due to poor coverage, DR effectively reduces to the Direct Method, which achieves the same consistency without the variance penalty from unstable weights (see Section 5, Q3). (iii) SIMCal-W convergence. Theorem 4 assumes SIMCal-W converges to $m^{\star}_{\rho}$; this holds under standard regularity conditions for isotonic regression and cross-validated stacking (see, e.g., vanderLaan2007SuperLearner).
5 Experiments
We evaluate CJE on a large-scale benchmark. We initially expected doubly robust methods to dominate by combining logged data with fresh draws. Instead, we found two surprises: (i) OPE fails catastrophically even after weight stabilization boosts ESS above 90%, and (ii) DR provides no advantage over Direct under low overlap, effectively reducing to the outcome model alone. These findings motivated the CLE diagnostic, which explains why high ESS is insufficient.
Our experiments address three questions:
  1. 1.
    Does calibration work? AutoCal-R should reduce RMSE, improve ranking accuracy, and with OUA dramatically improve CI coverage.
  2. 2.
    Why does OPE fail despite high ESS? We expected IPS to work after stabilization; CLE predicts failure when TTC is low.
  3. 3.
    Why doesn’t DR dominate? Under low overlap, DR’s IPS component contributes noise rather than information.
5.1 Experimental design
Data.
We started with 5,000 prompts from Chatbot Arena (Zheng2023LLMasJudge), a crowdsourced platform where users chat with anonymous LLMs. Prompts were randomly selected from first-turn English conversations; after filtering for teacher-forcing reliability, n=4,961n{=}4{,}961 samples remain with complete data for all policies (see Limitations, Section 6).
Policies.
Five LLM configurations, each generating responses to all prompts:
  • base: Llama 3.3 70B with standard system prompt (logging policy $\pi_{0}$)
  • clone: Same model and prompt as base, different seed (TF reliability test)
  • premium: Llama 3.1 405B (model size effect)
  • parallel_universe_prompt: Llama 3.3 70B with modified system prompt (prompt engineering test)
  • unhelpful: Deliberately low-quality, confusing responses (adversarial stress test)
Total: ${\sim}25$k responses ($\sim$}5k prompts $\times$ 5 policies; n=4,961n{=}4{,}961 after filtering).
Oracle and judge.
Oracle $Y$: GPT-5 (OpenAI2025GPT5) quality scores (0–100, normalized to $[0,1]$) collected for every response. Judge $S$: GPT-4.1-nano scores on all responses (temperature 0); approximately $16\times$ cheaper than oracle. Note: GPT-5 operates at temperature 1.0 due to API constraints, introducing stochasticity in oracle labels; we did not repeat-score responses, so coverage results may be conservative. All estimators except naive-direct calibrate judge scores using AutoCal-R. Base variants use monotone-only calibration (Z=SZ{=}S); +cov variants use two-stage calibration with response length as a covariate ($Z{=}g(S,\text{length})$). For DR methods, rewards $R=\hat{f}_{\mathrm{all}}(Z)$ use the full calibrator while the outcome model $\hat{q}=\hat{f}^{(-k)}(Z)$ uses cross-fitted (leave-one-fold-out) predictions; the difference $R-\hat{q}$ captures residual information for the IPS correction.
Experimental grid.
We cross 13 estimators $\times$ 5 oracle fractions (5%, 10%, 25%, 50%, 100%) $\times$ 5 sample sizes (500, 1k, 2k, 3k, 5k) $\times$ 50 random seeds = 16,250 total runs (1,250 per estimator). Reported metrics are means across seeds within each (oracle fraction, sample size) cell.
Metrics.
Ranking: Pairwise accuracy ($\text{PairAcc}=\binom{P}{2}^{-1}\sum_{j<k}\mathbf{1}\{(\hat{V}_{j}-\hat{V}_{k})(V^{\star}_{j}-V^{\star}_{k})>0\}$, the fraction of correctly ordered policy pairs), Top-1 accuracy, Kendall’s $\tau$. Magnitude: RMSEd (oracle-noise-debiased): $\mathrm{RMSE}^{d}=\sqrt{\max(0,\,\mathrm{MSE}-\widehat{\operatorname{Var}}(\bar{Y}))}$, where $\widehat{\operatorname{Var}}(\bar{Y})$ is the variance of the oracle mean estimate (conservatively bounded by $0.25/n$ per policy). This isolates estimator error from irreducible oracle sampling noise. Uncertainty: Coverage (95% CI capture rate). Note: Accuracy and uncertainty metrics exclude the unhelpful policy (which intentionally violates transportability); ranking metrics include all five policies to test whether estimators correctly identify the adversarial policy as worst.
5.2 Results
Table 1: Estimator comparison across all regimes. Metrics averaged over 5 sample sizes $\times$ 5 oracle fractions $\times$ 50 seeds. RMSEd = oracle-noise-debiased; CI coverage target = 95%.
Estimator RMSEd $\downarrow$ Coverage % SE $\downarrow$ Pairwise % $\uparrow$ Top-1 % $\uparrow$ $\tau$ $\uparrow$ Time (s)
Direct Methods (on-policy, no importance weights)
naive-direct 0.083 0.0 0.004 90.9 79.6 0.82 0.5
direct 0.023 85.7 0.007 91.8 84.1 0.84 0.9
direct+cov 0.026 87.0 0.008 94.3 89.4 0.89 2.9
IPS Methods (off-policy via importance weighting)
SNIPS 0.160 98.3 0.182 38.3 8.7 -0.24 0.3
SNIPS+cov 0.184 97.8 0.184 37.4 6.4 -0.25 0.5
calibrated-ips 0.025 99.1 0.096 47.1 19.1 -0.06 0.5
calibrated-ips+cov 0.026 99.2 0.103 46.5 20.6 -0.07 0.7
DR Methods (doubly robust, combines IPS + outcome model)
dr-cpo 0.046 99.2 0.037 78.4 46.3 0.57 6.5
dr-cpo+cov 0.057 99.4 0.045 79.1 50.1 0.58 20.3
calibrated-dr-cpo 0.023 99.5 0.023 91.0 81.0 0.82 6.6
calibrated-dr-cpo+cov 0.026 99.3 0.024 94.1 87.5 0.88 20.2
stacked-dr 0.023 95.5 0.013 91.9 83.1 0.84 14.6
stacked-dr+cov 0.028 96.1 0.013 92.1 84.7 0.84 62.5
$\downarrow$: lower is better; $\uparrow$: higher is better. Bold: best; underline: second-best. Coverage: closest to 95% is best.
Naming: All estimators except naive-direct use reward calibration (AutoCal-R). The “calibrated-” prefix adds weight stabilization (SIMCal-W). The “+cov” suffix uses two-stage calibration with response length.
Q1: Does calibration work?
Yes, for both accuracy and inference. Table 1 shows calibration improves accuracy across method families. For Direct: uncalibrated naive-direct achieves 90.9% pairwise vs. calibrated direct+cov at 94.3%. For DR: raw-weight dr-cpo achieves 78.4% pairwise ($\tau{=}0.57$) vs. weight-stabilized calibrated-dr-cpo at 91.0% ($\tau{=}0.82$), a 12.6 percentage point gain (both use AutoCal-R; the gain is from SIMCal-W). For inference: uncalibrated CIs (naive-direct) achieve 0% coverage—nominal 95% CIs capture the true value in 0/50 seeds. Calibration with OUA dramatically improves coverage: Direct achieves 85–87%, stacked-DR achieves near-nominal 95–96%. The OUA share of total variance ranges from 0% (100% oracle) to over 50% (5% oracle), explaining why ignoring calibration uncertainty produces catastrophically narrow CIs (Figure 5).
Q2: Why does OPE fail despite high ESS?
[…]
Q3: Why doesn’t DR dominate?
We expected DR to outperform Direct by combining logged data with fresh draws. Instead, direct+cov slightly outperforms calibrated-dr-cpo+cov (94.3% vs. 94.1%). The explanation: under low overlap, DR’s IPS component contributes noise rather than information. DR estimators are consistent if either the propensity model or the outcome model is correct. Given IPS’s failure, DR effectively relies entirely on the outcome model $\hat{q}(Z)$, not the propensities. This is exactly the regime CLE identifies: TTC is low, so logged data contributes little; DR $\approx$ Direct ++ extra variance from unstable weights. Implication: When teacher forcing is unreliable, prefer Direct over DR; it achieves the same accuracy without requiring logprobs.
Table 2: Ablation: marginal contribution of CJE components. Metrics averaged over 5 sample sizes $\times$ 5 oracle fractions $\times$ 50 seeds. CI coverage = % of 95% CIs containing truth.
[…]
Effect of AutoCal-R (Reward Calibration)
Direct, no calibration 0.083 0%
Direct + AutoCal-R 0.023 86% –72%
Effect of SIMCal-W (Weight Stabilization)
SNIPS (raw weights) 0.160 98%
Calibrated-IPS (+ SIMCal) 0.025 99% –84%
Combined Effect on DR
DR-CPO, raw weights 0.046 99%
DR-CPO + SIMCal 0.023 100% –50%
DR-CPO + SIMCal + cov 0.026 99% –43%
Recommended Configuration
Direct + AutoCal-R + cov 0.026 87% –69%
Key findings: AutoCal-R reduces RMSE by 72% and restores coverage (0%→86%). SIMCal-W reduces IPS RMSE by 84%. Both contribute additively.
Component ablation.
Table 2 isolates the marginal contribution of each CJE component. AutoCal-R (reward calibration) reduces RMSE by 72% and is necessary to avoid catastrophic undercoverage: without it, 0% of CIs contain the true value; with it, coverage improves to 86%. SIMCal-W (weight stabilization) reduces IPS RMSE by 84% while maintaining coverage, validating the $S$-monotone projection approach. For ranking accuracy, we recommend direct+cov (Direct + AutoCal-R with response-length covariate); for CI-critical applications, stacked-DR provides near-nominal coverage.
Table 3: Weight stabilization effect and CLE diagnostics by policy. SIMCal-W dramatically improves ESS while TTC reveals why IPS still fails: the logger rarely visits target-typical regions.
ESS (%) CLE Diagnostics
Policy SNIPS SIMCal Uplift TTC ($\hat{\beta}$) CLE Factor
[…]
TTC = Target-Typicality Coverage; CLE Factor = precision floor inflation. TTC $70\%$: logs-only IPS expected to fail regardless of ESS. Bold: passes TTC threshold.
A/A note: Clone’s raw ESS (26%) is below the theoretical 100%, reflecting API-level non-determinism and floating-point divergence in log-probability summation; SIMCal-W corrects this measurement artifact, restoring ESS to 99%. The 26% is still substantially higher than other policies (0.4–0.7%), confirming TF captures real distributional similarity. The dominant issue is CLE: LLM response trajectories are high-dimensional enough that even moderate policy differences yield poor coverage, making IPS unreliable regardless of TF quality. TTC (79.6% for clone vs. 19–49% for others) provides a more robust overlap diagnostic.
Weight diagnostics.
[…]
Figure 4: CJE output: policy value estimates at n=1000n{=}1000, 25% oracle. Red diamonds show oracle ground truth; blue circles show CJE estimates with 95% CIs. CIs capture the true value for policies satisfying transportability; unhelpful (which violates transportability) shows slight miscoverage, a failure mode flagged by the transportability test (Figure 6). This is representative of what a practitioner would see when applying CJE to their own evaluation data.
Figure 5: Uncertainty decomposition at n=1000n{=}1000 for direct+cov. Yellow: base sampling uncertainty (constant $\sim$0.014). Orange: oracle uncertainty (dominates at 5% oracle fraction, vanishes at 100%). At 5% oracle fraction, OUA contributes 55% of total variance; ignoring it produces invalid CIs.
Figure 6: Policy-wise mean transport test at 25% oracle fraction. Testing $H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ per policy (Bonferroni-corrected $\alpha{=}0.0125$). Clone, Parallel Universe, and Premium pass (residuals centered at zero). Unhelpful fails (mean residual =0.31=-0.31): the surrogate overestimates oracle scores for adversarial responses by +0.31+0.31 in absolute level.
Calibration transportability.
Figure 6 applies the policy-wise mean transport test (3) to validate whether the base-trained calibration can be reused across target policies. We test $H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ for each policy with Bonferroni correction ($\alpha{=}0.0125$). Clone, Parallel Universe, and Premium all pass: their residual distributions are centered at zero, implying that Direct + calibrated surrogate remains mean-unbiased for these policies. Unhelpful fails dramatically (mean residual 0.31-0.31, $p<0.001$). By 3, this implies $V^{\mathrm{oracle}}_{\pi^{\prime}}-V^{\mathrm{sur}}_{\pi^{\prime}}\approx-0.31$: the surrogate systematically overestimates oracle quality for adversarial responses. This flags a failure of mean transportability: absolute value estimates for this policy are biased, triggering the REFUSE-LEVEL gate (App. D.6). Practical implication: This test decides whether a single calibration can be safely reused across policies; if a policy fails, either recalibrate with policy-specific oracle data or fall back to oracle-only evaluation for that cell.
Sample size planning.
Figure 7 shows MDE contours for the Direct Method: the smallest true policy difference detectable at 80% power, 5% two-sided significance. Key observation: for a fixed label budget $m=n\times\text{oracle fraction}$, larger $n$ with lower oracle fraction outperforms smaller $n$ with higher oracle fraction, suggesting spread labels across more samples as a practical rule. Practical guidance: For new domains, oversample the first batch (e.g., n=2,000n{=}2{,}000 with 25% oracle fraction) to generate a domain-specific MDE plot, then use it to select the minimal viable $(n,\text{oracle fraction})$ for future experiments.
Figure 7: Sample size planning for direct+cov. (A) RMSE by sample size and oracle fraction. (B) Standard errors including OUA. (C) MDE contours showing the smallest effect detectable at 80% power. Dashed lines show common detection thresholds; cells with MDE <0.02<0.02 enable detecting 2-point quality differences. Example: To detect a 3% improvement with 80% power requires $n{\geq}1{,}000$ with ${\geq}10\%$ oracle fraction.
Cost analysis.
[…]
Method Oracle Cost Judge Cost Total Pairwise %
Pure oracle (100%) \$55.00 \$55.00 100%
CJE (5% oracle) \$0.55 \$3.50 \$4.05 99%
Table 5: Estimator Selection Guide. Requirements and recommended use cases for each estimator family.
Estimator Fresh Logprobs Overlap When to Use
Direct Default for LLM eval
IPS/SNIPS Required Small policy changes only
Calibrated IPS Required + SIMCal-W stabilization
DR Helpful Logs + fresh draws available
Stacked-DR Helpful Production ensemble
5.3 Practical takeaways
  1. 1.
    Default to Direct + two-stage calibration. If you can generate fresh responses on a shared prompt set, use the Direct Method. It requires no overlap assumption and achieves the best ranking accuracy (Table 5).
  2. 2.
    Check TTC before using OPE. If TTC <0.7<0.7, logs-only IPS will fail. Use DR only when fresh draws are available to complement logged data.
  3. 3.
    Always use OUA inference. Standard CIs severely under-cover. The jackknife cost is negligible.
  4. 4.
    Use OUA share to guide resource allocation. The fraction $\widehat{\mathrm{Var}}_{\mathrm{cal}}/\widehat{\mathrm{Var}}_{\mathrm{total}}$ identifies bottlenecks: if OUA share $50\%$, collect more oracle labels; if OUA share $20\%$, collect more evaluation prompts (Figure 5).
  5. 5.
    Covariates matter. Response length as a calibration covariate (not a reweighting covariate) improves ranking across all methods.
  6. 6.
    Optimize the budget ratio. How many oracle labels do you need? The optimal ratio of labels ($m$) to scores ($n$) follows a square-root law based on costs and variances (Appendix F). Check the OUA share $\omega$ to equalize marginal utility: if $\frac{\omega}{1-\omega}>\frac{\mathrm{Spend}_{\mathrm{oracle}}}{\mathrm{Spend}_{\mathrm{surrogate}}}$, the experiment is under-investing in oracle labels.
6 Limitations
Oracle alignment (scope).
We assume the operational oracle $Y$ aligns with stakeholder values; oracle selection itself is a governance question outside the scope of this work (Assumption A0 in App. C.13).
Overlap (positivity).
IPS requires support overlap between $\pi_{0}$ and each $\pi^{\prime}$; DR benefits from overlap but remains consistent via its outcome model when overlap is poor; Direct requires no overlap (it generates fresh responses under $\pi^{\prime}$). When overlap is poor, raw ratios are heavy-tailed and uncertainty inflates. Mitigations: SIMCal–W reduces dispersion and raises $\mathrm{ESS}$; if tails persist we (i) gate on $\mathrm{ESS}$ and Hill indices, (ii) use overlap weighting or cohort restriction, and (iii) run an online check when $\widehat{\alpha}_{\text{Hill}}{<}1$ or single–row dominance persists (App. D).
Judge assumptions (surrogate validity).
AutoCal–R assumes mean sufficiency and monotonicity in $S$ (or a learned index). If strained, the two–stage fallback preserves mean honesty but targets $\mathbb{E}[f(S^{\pi^{\prime}})]$. Mitigations: surface reliability curves and regional residuals; when evidence is weak, label as surrogate–target, widen/refresh the oracle slice, and target labels where error concentrates.
Calibration failure modes.
AutoCal-R can fail in three distinct ways, each with observable signatures:
  1. 1.
    Policy-specific calibration: The $S\to Y$ mapping differs across policies. Signature: Within-policy rankings accurate, cross-policy comparisons fail. Test: Policy-wise mean transport test (3). If $\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]\neq 0$, surrogate values are biased for that policy.
  2. 2.
    Extrapolation beyond range: Target policy produces $S$ values outside the calibration range. Signature: Boundary flatness; OutOfRange $5\%$. Test: Coverage diagnostic (App. D).
  3. 3.
    Temporal drift: The $S\to Y$ mapping changes over time. Signature: Recent residuals larger than historical; calibration accuracy degrades. Test: Periodic mean transport test on fresh oracle batches.
Scope of the mean transport test.
The policy-wise mean transport test (3) is a necessary but not sufficient condition for full surrogate transportability (S2): it guarantees that the surrogate is unbiased for policy-wise mean values, but does not rule out conditional miscalibration within subgroups or at the tails of the score distribution. Mitigations: perform subgroup-specific transport tests when fairness concerns apply; monitor reliability diagrams for regional deviations; use richer moment conditions for stronger tests (see remark after 3).
Calibration coverage (identification).
If a $\pi^{\prime}$ pushes $S$ outside the labeled range, isotonic calibration flattens at the boundary and levels are not point–identified. Mitigations: flag Limited Calibration Support and set Refuse-Level (report rankings and partial–ID bounds) until targeted labels cover the uncovered $S$ region (App. D).
Approximate sufficiency (bias modulus).
When $\mathbb{E}[Y\mid X,A,S]\neq\mu(S)$, the residual $\Delta(X,A,S)$ induces bias proportional to calibration error. Mitigations: by Cauchy–Schwarz, $|\mathrm{Bias}|\leq\|m-W_{\pi^{\prime}}\|_{2}\,\|\Delta\|_{2}$; DbP shrinks $\|m-W_{\pi^{\prime}}\|_{2}$, so bias is second order when either calibration is tight or the violation small; we surface this via diagnostics and invoke Refuse-Level when unbounded.
Temporal dependence and logger drift.
Non–stationarity (launches, safety updates) can bias or widen intervals. Mitigations: report dependence–robust SEs (block/stationary bootstrap), shorten analysis windows, and monitor judge drift via rank–based/residual change detection with FDR control.
Oracle independence and leakage.
OUA assumes the oracle slice is i.i.d. with non-overlapping folds. Mitigations: reuse deterministic folds across modules, de–duplicate the slice, and periodically refresh it.
Selection and multiplicity.
Scanning many $\pi^{\prime}$ inflates winner’s curse. Mitigations: use FDR control (BH/BY), optionally an outer split for IF–Stack to reduce selection optimism, and emphasize pre–specified contrasts.
Teacher forcing and API drift.
Accurate propensities require deterministic, chat–native TF (stable tokenizer/template) with additivity/conditionality invariants; missing/invalid TF corrupts ratios. In our experiments, 39 of 5,000 prompts (0.8%) were filtered: 11 due to base-policy TF conformance failures, and 28 due to missing target-policy logprobs. While small, we report this explicitly: practitioners should expect some data loss from TF gaps, and filtering should always be logged for reproducibility. This also highlights the current state of TF support from LLM providers: major closed-source providers (OpenAI, Anthropic, Google) do not currently support teacher forcing for chat models, and among open-weight hosting providers, Fireworks AI was the only service we found with reliable TF support. Even for our TF reliability test (clone = base with different seed), raw ESS is 26% rather than the theoretical 100%, reflecting API-level non-determinism and floating-point divergence in log-probability summation; SIMCal-W restores this to 99% (Table 3). This 26% is still substantially higher than other policies (0.4–0.7%), confirming TF captures real distributional similarity. The dominant issue is CLE, not TF quality: LLM trajectories are high-dimensional enough that even moderate policy differences yield poor coverage (see Table 3). Mitigations: enforce schema/conformance checks, ledger failures, and treat results as conditional on TF quality (App. E).
Subgroups and fairness.
Calibration quality and ESS gains may differ across subgroups. Mitigations: provide subgroup diagnostics (ESS, reliability) and, when feasible, use subgroup–specific calibration/weights or constrained pooling.
Judge informativeness (garbling).
Coarser judges raise the surrogate information bound and widen CIs. Mitigations: prefer richer rubrics (multi–dimensional $S$ with stable aggregation) and validate with coarsening ablations (empirical Blackwell monotonicity).
Compute.
DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$; AutoCal–R refits for OUA add modest overhead. We amortize via TF caches, shared folds, and a small stacking library.
Ethics Statement. We analyze retrospective logs that may include sensitive content. Diagnostics/gates prevent overconfident claims under poor overlap/coverage and surface judge drift. When identification fails, we report rankings only (Refuse-Level) and recommend targeted labeling or online checks. Any deployment should assess subgroup reliability and adopt privacy safeguards for logs.
Reproducibility Statement. Exact prompts, schema, TF contract, pseudocode, and numerics appear in the appendices. Code and data are available at https://github.com/cimo-labs/cje-arena-experiments; the CJE library is at https://github.com/cimo-labs/cje.
7 Conclusion
Uncalibrated LLM-as-judge evaluation is the default practice, and it produces rankings that can be inverted, confidence intervals with near-zero coverage, and OPE estimates that fail silently. These are not edge cases; they are the norm when calibration is ignored.
We introduced CJE, an audit-ready framework that fixes all three failures, built around a single rule: Design-by-Projection (DbP). The principle is simple: encode justified assumptions as closed convex sets and project valid objects onto them. Projections onto subspaces (nuisance-orthogonal scores), monotone cones (mean-preserving reward/weight calibration), and simplices (variance-hedged stacking) preserve the estimand while weakly reducing variance.
Concretely, AutoCal-R learns a mean-preserving surrogate $R=f(Z(S))$; SIMCal-W stacks $S$-monotone candidates to produce unit-mean ratios with empirical $\mathrm{ESS}$ gains via OOF stacking and a variance guard; DR-CPO delivers $\sqrt{n}$ inference under cross-fitting; IF-Stack minimizes plug-in IF variance; and OUA adds calibration uncertainty to dramatically improve CI coverage.
Theoretically, the Efficiency via Model Restriction result, applying established semiparametric principles to surrogate-based evaluation, explains why encoding justified knowledge as model restrictions lowers the efficiency bound, which is attainable with projection-designed estimators and cross-fitting. Two design corollaries follow: (i) Blackwell–efficiency monotonicity: richer judges (finer $\sigma$-fields) strictly help; (ii) under the exact $\mathrm{IsoMeanOne}_{S}$ projection, isotonic mean-one calibration weakly reduces weight dispersion (the reference implementation uses a simpler normalization that empirically yields similar $\mathrm{ESS}$ gains). Empirically, on Arena-derived logs, SIMCal-W turns near-degenerate ratios into stable weights (large $\mathrm{ESS}$ gains), stacked-DR achieves near-nominal coverage and near-$\sqrt{n}$ scaling, and stacking improves ordering; when calibration support is limited, CJE flags the issue and reports robust rankings with conservative uncertainty (Refuse-Level).
Takeaways.
(i) Default to Direct: OPE fails under low overlap even with high ESS; calibrated judge scores suffice when transportability holds.
(ii) Always use OUA: naive CIs achieve 0% coverage; OUA improves to 85–96%.
(iii) Validate transportability: the mean transport test catches real failures (e.g., adversarial policies).
Future work.
Selection-aware inference over large policy sets; robust/DP isotonic calibration (mirror/Bregman DbP) for heavy tails; active oracle budgeting via shadow prices; sequential/agent evaluations with prefix-aware SIMCal and stepwise DR; and subgroup-aware constraints with fairness diagnostics.
8 Related Work
CJE draws on four research threads: surrogate-assisted causal inference, LLM evaluation, probability calibration, and off-policy evaluation. We position our contributions by first clarifying three regimes of surrogacy that organize the literature.
Three regimes of surrogacy.
Surrogate-assisted estimation admits three regimes with increasing assumptions and decreasing oracle burden:
  1. 1.
    Surrogates for efficiency only (KallusMao2020Surrogates): The surrogate $S$ improves estimation efficiency (e.g., as features in outcome models) but does not identify treatment effects. Oracle labels $Y$ are required in every evaluation context. This is the fallback when stronger surrogacy assumptions fail.
  2. 2.
    Local surrogacy (AtheyChettyImbensKang2019; Prentice1989Surrogate): Surrogate sufficiency (S1) holds within a single environment $g^{\star}$: $\mathbb{E}[Y\mid X,A,S]=f(S,X)$ in $g^{\star}$. Calibrate once per environment; evaluate many policies within that environment using only $S$.
  3. 3.
    Global surrogacy with transport (CJE): Surrogate sufficiency (S1) plus transportability (S2): the same calibration $f$ applies across environments. Calibrate once; evaluate policies across environments using only $S$, subject to transport diagnostics.
CJE operates in Regime 3. This requires stronger assumptions than Kallus & Mao (who use surrogates for efficiency but not identification) and differs in scope from Athey et al. (who build a surrogate index for binary-treatment ATEs in a single environment). The benefit is amortization: one oracle slice enables evaluation across policies and deployment contexts. The cost is that S2 must hold; our diagnostics (TTC, Bhattacharyya affinity, transport tests) validate when it does.
LLM-as-judge evaluation.
Automatic judges (LLMs scoring other LLMs) are now standard for scalable evaluation (Zheng2023LLMasJudge; Kim2024PrometheusJudge; KocmiFedermann2023). Benchmarks like AlpacaEval (DuboisLiTaoriEtAl2024AlpacaEval) and RewardBench (LambertPiechChenEtAl2024RewardBench) systematically compare judge quality, revealing biases (verbosity, position, self-preference) and drift under distribution shift (DietzEtAl2025LLMJudges). These works focus on ranking judges by correlation with human labels; CJE addresses the orthogonal problem of calibrating a fixed judge to produce unbiased point estimates and calibration-aware confidence intervals for policy value. Recent work like CalibraEval (Li2025CalibraEval) addresses position and token bias via label-free calibration; CJE is complementary, providing oracle-grounded calibration with transportability guarantees and uncertainty quantification.
Probability calibration.
Platt scaling (Platt1999), isotonic regression (ZadroznyElkan2002; NiculescuMizilCaruana2005), and temperature scaling (GuoPleissSunWeinberger2017Calibration) are standard post-hoc calibration methods for classification. Our AutoCal-R applies isotonic regression not for classification confidence but for reward calibration: mapping a continuous judge score $S$ to an expected oracle outcome $\mathbb{E}[Y\mid S]$. The mean-preserving property (Lemma 1) ensures the expected value of calibrated rewards matches the oracle under the logging distribution.
Off-policy evaluation (OPE).
IPS (HorvitzThompson1952; Hajek1964Rejective) and doubly robust estimators (BangRobins2005; JiangLi2016; KallusUehara2020DRL) are foundational. Weight stabilization via clipping (Ionides2008TruncatedIS), balancing (Kallus2018Balanced), and overlap weighting (LiMorganZaslavsky2018; Crump2009LimitedOverlap) address variance inflation. Most relevant is recent work on isotonic-calibrated IPW (vanDerLaanLinCaroneLuedtke2025StabIPW) and DR inference via calibration (vanDerLaanLuedtkeCarone2024DRCalibration). Our SIMCal-W extends this by projecting importance weights onto the cone of judge-score-monotone functions, exploiting surrogate structure absent in generic OPE.
Ensemble methods for estimation.
Super Learner (vanderLaan2007SuperLearner) provides a powerful framework for combining candidate algorithms by minimizing cross-validated prediction risk. Recent work extends ensemble ideas to semiparametric estimation (ShinLiuColeFine2020EnsembleSemiparametric), combining efficient estimators derived from factored likelihoods. Our IF-space stacking (Theorem 5) takes a complementary approach: instead of combining predictions to minimize MSE, we combine estimators by minimizing the asymptotic variance of their combined influence function. This provides a general formulation that directly targets efficiency, is not tied to specific likelihood factorizations, and is explicitly framed as optimization in the Hilbert space of influence functions.
Transportability.
Generalizing from source to target populations is studied in transportability (PearlBareinboim2014) and covariate shift (Shimodaira2000CovShift; SugiyamaKrauledatMueller2007IWCV). Our S2 assumption is a transportability condition: calibration learned on the oracle slice applies to the full dataset (and across policies) when no selection mechanism points into $Y$ given $(X,A,S)$. The CLE bound (Section 2) formalizes when logs-only estimation fails even with valid transportability, motivating the shift to Direct or DR methods under poor coverage.
Acknowledgements
We thank Kevin Zielnicki, Molly Davies, Sven Schmit, Brad Klingenberg, Chris Rinaldi, Viridiana Lourdes, Izzy Farley, Sudhish Kasaba Ramesh, Sean Kelly, and Adith Swaminathan for helpful discussions and feedback on earlier drafts of this work.
Appendix A Notation and Formal Setup
Observed data and policies.
We observe i.i.d. logs
$O_{i}=(X_{i},A_{i},S_{i},Y^{\mathrm{obs}}_{i},L_{i}),\qquad i=1,\dots,n,$
generated under a fixed logger $A_{i}\sim\pi_{0}(\cdot\mid X_{i})$. A scalar judge $S_{i}=s(X_{i},A_{i})$ is available on all rows. The label indicator $L_{i}\in\{0,1\}$ marks inclusion in the oracle slice; when $L_{i}=1$ we observe $Y^{\mathrm{obs}}_{i}=Y_{i}$, otherwise $Y^{\mathrm{obs}}_{i}$ is missing. For a candidate policy $\pi^{\prime}$, define the sequence-level importance ratio
$W_{\pi^{\prime},i}\;=\;\frac{\pi^{\prime}(A_{i}\mid X_{i})}{\pi_{0}(A_{i}\mid X_{i})}\;=\;\exp\!\Big\{\log p_{\pi^{\prime}}(A_{i}\!\mid\!X_{i})-\log p_{\pi_{0}}(A_{i}\!\mid\!X_{i})\Big\},$
computed via teacher forcing (TF) with the model’s own tokenizer/rendering. Write
$W^{\mathrm{m1}}_{\pi^{\prime},i}\;=\;\frac{W_{\pi^{\prime},i}}{\tfrac{1}{n}\sum_{j=1}^{n}W_{\pi^{\prime},j}}$
for the sample–mean–one (SNIPS) baseline (global normalization over the evaluation cohort).
Estimand.
Let $Y(\pi^{\prime})$ denote the outcome under the counterfactual draw $A\sim\pi^{\prime}(\cdot\mid X)$. The target is
$V(\pi^{\prime})=\mathbb{E}\!\big[Y(\pi^{\prime})\big].$
A.1 Assumptions (compact)
(D1) Fixed logger & i.i.d. $(X_{i},A_{i},S_{i})$ are i.i.d. under $\pi_{0}$; TF log-likelihoods are stable and well-defined.
(D2) Overlap (positivity). $\pi_{0}(a\mid x)>0$ whenever $\pi^{\prime}(a\mid x)>0$, and $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}]<\infty$.
(D3) Judge coverage & stability. $S$ is well-defined under both $\pi_{0}$ and $\pi^{\prime}$; the Radon–Nikodym derivative on $\sigma(S)$ exists; the judge/rubric is stable on the analysis window.
(J1) Oracle slice. There exists an i.i.d. subsample $O=\{i:L_{i}=1\}$ with $m=|O|\ll n$ on which $Y$ is observed.
(J2-M) Mean sufficiency (monotone). $\mathbb{E}[Y\mid X,A,S]=\mu(S)$ with $\mu$ weakly nondecreasing. (J2-SI) Single-index fallback. There exist $g^{\star}:\mathcal{S}\times\mathcal{X}\!\to\!\mathbb{R}$ and nondecreasing $\mu^{\star}$ such that $\mathbb{E}[Y\mid X,A,S]=\mu^{\star}\!\big(g^{\star}(S,X)\big)$. When $g^{\star}$ depends only on $S$, this reduces to J2-M; when covariates $X$ (e.g., response length) are included in $g$, the target remains $V(\pi^{\prime})$ and the estimand is preserved provided the index is consistently estimated.
(R1) Tails/moments. $\mathbb{E}[Y^{2}]<\infty$, $\mathbb{E}[S^{2}]<\infty$, and $\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}]<\infty$. Under the exact $\mathrm{IsoMeanOne}_{S}$ projection with relative variance bound $\rho\geq 1$, $\operatorname{Var}_{n}(\hat{W}_{\pi^{\prime}})\leq\rho\,\operatorname{Var}_{n}(W^{\mathrm{m1}}_{\pi^{\prime}})$ (i.e., variance at most $\rho$ times the SNIPS baseline).
(R2) Calibration consistency. AutoCal-R satisfies $\|\hat{f}(Z)-\mathbb{E}[Y\mid Z]\|_{L^{2}(P_{Z})}=o_{p}(1)$, where $Z=S$ in monotone mode and $Z=g(S,X)$ (learned index) in two-stage mode. SIMCal-W satisfies $\|\hat{W}_{\pi^{\prime}}-m^{\star}_{\rho}\|_{L^{2}(P)}=o_{p}(1)$, where $m^{\star}_{\rho}$ is the $S$-monotone, mean-one, variance-capped projection of $W_{\pi^{\prime}}$ (see Section A.3); when $\mathbb{E}[W_{\pi^{\prime}}\mid S]$ is itself monotone and satisfies the variance constraint, $m^{\star}_{\rho}=\mathbb{E}[W_{\pi^{\prime}}\mid S]$.
(R3) One-of-two rates with cross-fitting. With nuisances $\hat{q}(x,a)\approx\mathbb{E}[R\mid x,a]$ and $\hat{W}_{\pi^{\prime}}$,
$\|\hat{q}-q^{\star}_{R}\|_{L^{2}(P)}\cdot\|\hat{W}_{\pi^{\prime}}-m^{\star}\|_{L^{2}(P)}\;=\;o_{p}(n^{-1/2}),$
e.g., either factor is $o_{p}(n^{-1/4})$ with the other consistent.
A.2 Cross-fitting and folds
Let $F:\{1,\dots,n\}\!\to\!\{1,\dots,K\}$ be a deterministic fold map (e.g., a hash of x_idx\_id). For any learner $\mathcal{L}$, train $\hat{\eta}^{(-k)}=\mathcal{L}$ on $\{i:F(i)\neq k\}$ and use out-of-fold predictions $\hat{\eta}^{\mathrm{OOF}}_{i}=\hat{\eta}^{(-F(i))}(O_{i})$ in influence-function (IF) calculations. The same folds are reused across AutoCal-R, SIMCal-W, and DR nuisances.
A.3 Projection operators used by CJE
Monotone cone. $\mathcal{M}_{\uparrow}=\{f:\mathbb{R}\!\to\!\mathbb{R}\text{ nondecreasing}\}$. The isotonic projector (PAVA) $\Pi_{\mathcal{M}_{\uparrow}}$ enjoys: (i) $L^{2}$ optimality; (ii) mean preservation on the training sample; (iii) dispersion reduction by majorization.
Mean-one cone for weights. For $w\in\mathbb{R}^{n}$ ordered by $S$, define the mean-one isotonic projection
$\mathrm{IsoMeanOne}_{S}(w)\;=\;\mathop{\mathrm{arg\,min}}\limits_{u}\ \sum_{i}(u_{i}-w_{i})^{2}\quad\text{s.t.}\quad u\in\mathcal{M}_{\uparrow}(S),\ \tfrac{1}{n}\sum_{i}u_{i}=1,$
where $\mathcal{M}_{\uparrow}(S)$ denotes vectors nondecreasing in the $S$-order. This preserves the sample mean and weakly reduces empirical variance; hence $\mathrm{ESS}$ weakly increases (deterministically, by majorization).
Simplex hull. For centered IF columns $\{\phi^{(e)}\}_{e\in\mathcal{E}}$, let $\Phi=[\phi^{(e)}]$ and $\Delta=\{\alpha:\alpha_{e}\geq 0,\ \sum_{e}\alpha_{e}=1\}$. IF-space stacking solves $\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$ with $\hat{\Sigma}=(1/n)\Phi^{\top}\Phi+\lambda I$.
A.4 AutoCal-R and SIMCal-W primitives
AutoCal-R. On $O=\{i:L_{i}=1\}$, fit $R=\hat{f}(Z(S))$ by either: (i) monotone ($Z(S)=S$), or (ii) two-stage ($Z(S,X)=\mathrm{ECDF}\{g(S,X)\}$ with a spline $g$), selecting by OOF RMSE (1-SE rule). Use OOF predictions $R^{\mathrm{OOF}}$ along IF paths; the point estimate may use a pooled fit. A terminal isotonic step enforces slice-mean preservation.
Figure 8: Two-stage calibration with covariates. Left: first-stage spline on judge score (response length fixed at mean). Center: first-stage effect of response length (judge score fixed). Right: second-stage isotonic mapping on the risk index (Stage 1 output) enforces monotonicity while preserving flexibility from the first stage.
SIMCal-W. (Per fold) Fit up/down isotonic maps on $S$ to obtain $W^{\mathrm{OOF}}_{\uparrow},W^{\mathrm{OOF}}_{\downarrow}$; optionally include $W^{\mathrm{OOF}}_{\mathrm{base}}\equiv 1$. Define residuals $\Delta_{i}$ (IPS: $R_{i}$; DR: $R_{i}-\hat{q}(X_{i},A_{i})$). Choose $\hat{\beta}\in\arg\min_{\beta\in\Delta_{3}}\beta^{\top}\hat{\Sigma}\,\beta$ where $\hat{\Sigma}$ is the covariance of $U_{c}=W^{\mathrm{OOF}}_{c}\Delta$. Form $W^{\mathrm{stack}}=\sum_{c}\hat{\beta}_{c}W^{\mathrm{OOF}}_{c}$, renormalize to mean one (global, over the evaluation cohort), optionally apply the absolute variance guard
$\alpha=\min\!\Big\{1,\ \sqrt{\frac{\rho}{\operatorname{Var}(W^{\mathrm{stack}})}}\Big\},\qquad W^{\mathrm{blend}}=1+\alpha\,(W^{\mathrm{stack}}-1),$
and normalize to mean one: $\hat{W}_{\pi^{\prime}}=W^{\mathrm{blend}}/\bar{W}^{\mathrm{blend}}$. (The theoretical Lemma 5 uses a relative bound.)
A.5 DR nuisances and sequence value
Let $\hat{q}(x,a)\approx\mathbb{E}[R\mid x,a]$ and define $\hat{g}_{\pi^{\prime}}(x)=\sum_{a}\pi^{\prime}(a\mid x)\,\hat{q}(x,a)$. For sequences, approximate $\hat{g}_{\pi^{\prime}}(x)$ with one (default) rollout $A^{\prime}\!\sim\!\pi^{\prime}(\cdot\mid x)$ and $R^{\prime}=\hat{f}\!\big(s(x,A^{\prime})\big)$; a light smoother (e.g., ridge over $(x,z)$ features) can reduce Monte Carlo noise. Cross-fitting is used throughout.
A.6 Influence functions and variance
Let $\{\phi_{i}\}_{i=1}^{n}$ denote the (approximately) centered influence–function contributions of $\hat{\psi}$, computed with cross–fitted/OOF nuisances and $R^{\mathrm{OOF}}$ along the IF path, so that $\frac{1}{n}\sum_{i=1}^{n}\phi_{i}\approx 0$. Under standard regularity conditions,
$\sqrt{n}\,\bigl(\hat{\psi}-\psi\bigr)\ \overset{d}{\longrightarrow}\ \mathcal{N}\!\left(0,\ \operatorname{Var}(\phi)\right),\quad\text{hence}\quad\operatorname{Var}(\hat{\psi})\approx\frac{\operatorname{Var}(\phi)}{n}.$
Since $\mathbb{E}[\phi]=0$, we have $\operatorname{Var}(\phi)\approx\frac{1}{n}\sum_{i=1}^{n}\phi_{i}^{2}$. We estimate the variance of $\hat{\psi}$ (not $\phi$) and the total variance (including the oracle addition) by
$\widehat{\operatorname{Var}}_{\mathrm{main}}\;=\;\frac{1}{n^{2}}\sum_{i=1}^{n}\phi_{i}^{2},\qquad\widehat{\operatorname{Var}}_{\mathrm{total}}\;=\;\widehat{\operatorname{Var}}_{\mathrm{main}}+\widehat{\operatorname{Var}}_{\mathrm{cal}},$
and report the $(1-\alpha)$ Wald interval
$\mathrm{CI}_{1-\alpha}:\ \hat{\psi}\ \pm\ z_{1-\alpha/2}\,\sqrt{\widehat{\operatorname{Var}}_{\mathrm{total}}}\,.$
When serial or cluster dependence is a concern, we additionally report dependence–robust SEs (e.g., cluster–robust sandwich or block/stationary bootstrap) as a sensitivity analysis.
A.7 Oracle–uncertainty–aware (OUA) jackknife
Partition $O$ into $K$ oracle folds $\{O_{k}\}_{k=1}^{K}$. For each $k$, refit AutoCal-R on $O\setminus O_{k}$, recompute $R^{(-k)}$, and rerun the full pipeline to obtain $\hat{\psi}^{(-k)}$. Then
$\bar{\psi}=\tfrac{1}{K}\sum_{k=1}^{K}\hat{\psi}^{(-k)},\qquad\widehat{\operatorname{Var}}_{\mathrm{cal}}=\frac{K-1}{K}\sum_{k=1}^{K}\big(\hat{\psi}^{(-k)}-\bar{\psi}\big)^{2}.$
(With unequal fold sizes, use the standard weighted delete-one-group formula.)
A.8 Diagnostics (definitions)
ESS. $\mathrm{ESS}(W)=\big(\sum_{i}W_{i}\big)^{2}/\sum_{i}W_{i}^{2}$; we report the fraction $\mathrm{ESS}/n$. Under global mean-one normalization (SNIPS), $\sum_{i}W_{i}=n$ and
$\frac{\mathrm{ESS}(W)}{n}\;=\;\frac{1}{1+\mathrm{CV}^{2}(W)}\quad\text{with}\quad\mathrm{CV}^{2}(W)=\operatorname{Var}(W)\ \text{(since }\mathbb{E}[W]=1\text{)}.$
Unless stated otherwise, diagnostics use the global (not per-fold) mean-one scaling.
Max-weight share. $\max_{i}W_{i}\big/\sum_{j}W_{j}$.
Tail index (Hill). For top-$k$ order statistics $W_{(1)}\geq\cdots\geq W_{(k)}$, $\hat{\alpha}^{-1}=\tfrac{1}{k}\sum_{j=1}^{k}\log\!\big(W_{(j)}/W_{(k)}\big)$ (we sweep $k$ over a stability grid and report the plateau).
Bhattacharyya affinity in $S$. $A_{B}=\int\sqrt{p_{S|\pi^{\prime}}(s)\,p_{S|\pi_{0}}(s)}\,ds$ (discrete: sum over bins); $D_{B}=-\log A_{B}$.
DR orthogonality score. $n^{-1}\sum_{i}\hat{W}_{\pi^{\prime},i}\big(R^{\mathrm{OOF}}_{i}-\hat{q}^{\mathrm{OOF}}(X_{i},A_{i})\big)$ with a Wald CI.
Coverage badge. Plug-in estimate of $\Pr_{\pi^{\prime}}\!\big(S\notin[S_{\min}^{\mathrm{orc}},S_{\max}^{\mathrm{orc}}]\big)$; large out-of-range mass with near-flat boundaries triggers Limited Calibration Support and Refuse-Level.
A.9 Symbol glossary
Symbol Meaning
$X,A$ Context, action (sequence)
$S=s(X,A)$ Judge score (scalar)
$Y$ Ground-truth outcome (on oracle slice)
$\pi_{0},\pi^{\prime}$ Logger and candidate policies
$W_{\pi^{\prime}}$ Importance ratio $\pi^{\prime}(A\mid X)/\pi_{0}(A\mid X)$
$W^{\mathrm{m1}}_{\pi^{\prime}}$ Mean-one (SNIPS) baseline
$R=\hat{f}(Z(S))$ Calibrated reward (AutoCal-R; monotone or two-stage)
$\hat{W}_{\pi^{\prime}}$ Calibrated, unit-mean, $S$-monotone weights (SIMCal-W)
$\hat{q},\ \hat{g}_{\pi^{\prime}}$ Outcome and policy–value nuisances for DR
$\phi$ Per-row centered influence-function contribution
$\widehat{\operatorname{Var}}_{\mathrm{main}}$ Estimator variance $n^{-2}\sum_{i}\phi_{i}^{2}$
$\widehat{\operatorname{Var}}_{\mathrm{cal}}$ Calibration variance (oracle jackknife)
$\mathrm{ESS}(W)$ Effective sample size
Appendix B Algorithms (extended)
This appendix gives compact, cross-fitted pseudocode for CJE modules: reward calibration (AutoCal-R), surrogate-indexed weight calibration (SIMCal-W), estimators (Cal-IPS, DR-CPO), IF-space stacking, and oracle–uncertainty–aware variance (OUA). We reuse the same $K$-fold map $F(i)\!\in\!\{1{:}K\}$ across all modules. “OOF” denotes out-of-fold predictions used along the IF path.
Algorithm 1 AutoCal-R: mean-preserving reward calibration (cross-fitted; automatic two-stage fallback)
1:Inputs: Oracle pairs $\{(S_{i},Y_{i}):L_{i}{=}1\}$; optional covariates $\{X_{i}\}$; folds $F(\cdot)$; smooth index class $g(\cdot)$
2:Outputs: Calibrated rewards $R_{i}$ and OOF $R^{\mathrm{OOF}}_{i}$
3:for $k=1$ to $K$ do
4:  Train set $O_{\neg k}\!=\!\{i:L_{i}{=}1,\,F(i)\neq k\}$; test set $O_{k}\!=\!\{i:L_{i}{=}1,\,F(i)=k\}$
5:  Monotone candidate: $\hat{f}^{(-k)}_{\uparrow}\!\in\!\arg\min_{f\in\mathcal{M}_{\uparrow}}\sum_{i\in O_{\neg k}}(Y_{i}-f(S_{i}))^{2}$; set $R^{\mathrm{OOF}}_{\uparrow,i}\!=\!\hat{f}^{(-k)}_{\uparrow}(S_{i})$ for $i\!\in\!O_{k}$
6:  Two-stage candidate: fit $g^{(-k)}(S,X)$ on $O_{\neg k}$; ranks $U_{i}\!=\!\mathrm{ECDF}_{O_{\neg k}}(g^{(-k)}(S_{i},X_{i}))$; fit $\hat{h}^{(-k)}_{\uparrow}\!\in\!\arg\min_{h\in\mathcal{M}_{\uparrow}}\sum_{i\in O_{\neg k}}(Y_{i}-h(U_{i}))^{2}$; set $R^{\mathrm{OOF}}_{2s,i}\!=\!\hat{h}^{(-k)}_{\uparrow}(U_{i})$ for $i\!\in\!O_{k}$
7:end for
8: Compute OOF risks (overall and by $S$ tertile); select mode via 1-SE rule (prefer monotone unless two-stage is significantly better)
9: Refit the selected mode on the full oracle slice to obtain global $R_{i}$ for all $i\in\{1{:}n\}$; retain $R^{\mathrm{OOF}}_{i}$ per fold for IFs
10:Note: The terminal isotonic step preserves the oracle-slice mean exactly.
Algorithm 2 SIMCal-W: surrogate-indexed, unit-mean monotone weight calibration (OOF project $\to$ stack $\to$ cap $\to$ normalize)
1:Inputs: Baseline mean-one ratios $W^{\mathrm{m1}}_{\pi^{\prime}}$; scores $S$; residuals $\Delta$ (IPS: $\Delta{=}R$; DR: $\Delta{=}R-\hat{q}$); folds $F(\cdot)$; variance cap $\rho\!\geq\!1$ (default 11)
2:Output: Calibrated weights $\hat{W}_{\pi^{\prime}}$ (mean-one, approximately $S$-monotone)
3:for $k=1$ to $K$ do
4:  Train $I_{\neg k}=\{i:F(i)\neq k\}$; test $I_{k}=\{i:F(i)=k\}$ {OOF candidate projections}
5:  Fit isotonic regression of $W^{\mathrm{m1}}_{\pi^{\prime}}$ on $S$ over $I_{\neg k}$: increasing $m^{(-k)}_{\uparrow}(S)$ and decreasing $m^{(-k)}_{\downarrow}(S)$ (via S-S); rescale each to mean one on $I_{\neg k}$
6:  Predict on $I_{k}$: $W^{\mathrm{OOF}}_{\uparrow}=m^{(-k)}_{\uparrow}(S)$, $W^{\mathrm{OOF}}_{\downarrow}=m^{(-k)}_{\downarrow}(S)$, and optionally include $W^{\mathrm{OOF}}_{\mathrm{base}}\equiv 1$
7:end for
8:OOF stacking (variance-aware). Form $U_{c}=W^{\mathrm{OOF}}_{c}\,\Delta$ for $c\in\{\mathrm{base},\uparrow,\downarrow\}$; compute $\hat{\Sigma}_{cd}=\mathrm{cov}(U_{c},U_{d})+\lambda\mathbf{1}_{c=d}$
9: Choose $\hat{\beta}\in\arg\min_{\beta\in\Delta_{3}}\beta^{\top}\hat{\Sigma}\,\beta$; set $W^{\mathrm{stack}}=\sum_{c}\hat{\beta}_{c}\,W^{\mathrm{OOF}}_{c}$; renormalize $W^{\mathrm{stack}}$ to sample mean one
10:Light variance guard (optional). $\alpha=\min\!\left\{1,\ \sqrt{\rho/\operatorname{Var}(W^{\mathrm{stack}})}\right\}$; set $W^{\mathrm{blend}}=1+\alpha\,(W^{\mathrm{stack}}-1)$
11:Mean normalization. $\hat{W}_{\pi^{\prime}}\leftarrow W^{\mathrm{blend}}/\bar{W}^{\mathrm{blend}}$ (restores sample mean one)
Algorithm 3 Cal-IPS: calibrated importance sampling
1:Inputs: Calibrated rewards $R,R^{\mathrm{OOF}}$ (Alg. 1); calibrated weights $\hat{W}_{\pi^{\prime}}$ (Alg. 2)
2:Outputs: $\widehat{V}_{\mathrm{IPS}}$ and IF $\phi^{\mathrm{IPS}}$
[…]
Algorithm 4 DR-CPO: sequence-aware doubly robust estimator (cross-fitted)
1:Inputs: $\hat{W}_{\pi^{\prime}}$; $R$ and $R^{\mathrm{OOF}}$ from Alg. 1; folds $F(\cdot)$
2:Outputs: $\widehat{V}_{\mathrm{DR}}$ and IF $\phi^{\mathrm{DR}}$
3:for $k=1$ to $K$ do
4:  Train $\hat{q}^{(-k)}(x,a)\!\approx\!\mathbb{E}[R\mid x,a]$ on $\{i:F(i)\neq k\}$; predict OOF $\hat{q}^{\mathrm{OOF}}_{i}=\hat{q}^{(-k)}(X_{i},A_{i})$ for $i\in I_{k}$
5:  Approximate $\hat{g}^{(-k)}_{\pi^{\prime}}(x)=\mathbb{E}_{A\sim\pi^{\prime}(\cdot\mid x)}[\hat{q}^{(-k)}(x,A)]$ via one rollout $A^{\prime}\sim\pi^{\prime}(\cdot\mid X)$ and optional smoothing; obtain OOF $g^{\mathrm{OOF}}_{\pi^{\prime},i}$ for $i\in I_{k}$
6:end for
[…]
Algorithm 5 IF-Stack: variance-optimal convex ensembling in IF space
1:Inputs: Candidates $\{\widehat{V}^{(e)},\phi^{(e)}\}_{e\in\mathcal{E}}$ (centered IFs; same folds); ridge $\lambda$
2:Outputs: $\widehat{V}^{(\hat{\alpha})}$ and $\phi^{(\hat{\alpha})}$
3: Form $\Phi=[\phi^{(e)}]_{e\in\mathcal{E}}$; $\hat{\Sigma}=(1/n)\Phi^{\top}\Phi+\lambda I$
4: Solve $\hat{\alpha}\in\arg\min_{\alpha\in\Delta}\alpha^{\top}\hat{\Sigma}\,\alpha$
[…]
6:Optional outer split: learn $\hat{\alpha}$ on one half; apply to the other to reduce selection optimism
7:Support note (Carathéodory). If $\mathrm{rank}(\Phi)=r$, the variance-optimal stack uses at most r+1r{+}1 base estimators.
Algorithm 6 OUA jackknife: oracle–uncertainty–aware variance addition
[…]
2:Outputs: $\widehat{\operatorname{Var}}_{\mathrm{cal}}$ and $\widehat{\operatorname{Var}}_{\mathrm{total}}=\widehat{\operatorname{Var}}_{\mathrm{main}}+\widehat{\operatorname{Var}}_{\mathrm{cal}}$
3:for $k=1$ to $K$ do
4:  Refit AutoCal-R on $O\setminus O_{k}$; recompute $R^{(-k)}$ and all downstream nuisances & weights; run the full pipeline to get $\widehat{V}^{(-k)}$
5:end for
[…]
7:Return: $\widehat{\operatorname{Var}}_{\mathrm{total}}=\widehat{\operatorname{Var}}_{\mathrm{main}}+\widehat{\operatorname{Var}}_{\mathrm{cal}}$
Complexity notes.
PAVA is $O(n)$ after a shared sort by $S$. SIMCal-W is linear-time per fold; the stacking QP is $3{\times}3$ (weights) or a small $|\mathcal{E}|{\times}|\mathcal{E}|$ system. DR-CPO adds one rollout + judge per $(X,\pi^{\prime})$. The OUA jackknife refits the calibrator $K$ times and reruns the pipeline; TF caches and precomputed features amortize cost.
Appendix C Proofs and Technical Lemmas
We collect standing identities, shape-constrained facts, and proofs for the results in Section 4. Unless stated otherwise, expectations are under the logging law $P_{\pi_{0}}$; $L^{2}$ norms are with respect to the relevant marginal (e.g., $L^{2}(P)$ or $L^{2}(P_{S})$). We reuse the fold map $F(\cdot)$ from Section A.2 and the projection operators from Section A.3.
C.1 Standing identities and tools
Change of measure.
For any integrable $h(X,A,S,Y)$ and any candidate $\pi^{\prime}$,
$\mathbb{E}\!\big[W_{\pi^{\prime}}\,h(X,A,S,Y)\big]\;=\;\mathbb{E}_{\pi^{\prime}}\!\big[h(X,A,S,Y)\big],\qquad\mathbb{E}[W_{\pi^{\prime}}]=1.$ (2)
Doob–Dynkin / conditional expectation as $L^{2}$ projection.
Let $\mathcal{G}=\sigma(S)$. Then $m^{\star}(S):=\mathbb{E}[W_{\pi^{\prime}}\mid\mathcal{G}]$ is the $L^{2}$ projection of $W_{\pi^{\prime}}$ onto the closed subspace $L^{2}(\mathcal{G})\subset L^{2}(P)$, i.e.,
$\mathbb{E}\!\big[(W_{\pi^{\prime}}-U(S))^{2}\big]=\mathbb{E}\!\big[(W_{\pi^{\prime}}-m^{\star}(S))^{2}\big]+\mathbb{E}\!\big[(m^{\star}(S)-U(S))^{2}\big],$ (3)
for all $U(S)\in L^{2}(\mathcal{G})$. In particular, $\operatorname{Var}(W_{\pi^{\prime}}U)\geq\operatorname{Var}(m^{\star}(S)U)$ for any $U(S)\in L^{2}(\mathcal{G})$.
Pythagoras in Hilbert spaces.
Let $L^{2}_{0}(P)$ be the mean-zero Hilbert space with inner product $\langle f,g\rangle=\mathbb{E}[fg]$. For a nonempty closed convex set $\mathcal{C}\subset L^{2}_{0}$ and any $z\in L^{2}_{0}$, denote by $\Pi_{\mathcal{C}}(z)$ the metric projection. If $\mathcal{C}_{1}\subseteq\mathcal{C}_{2}$ then $\mathrm{dist}(z,\mathcal{C}_{2})\leq\mathrm{dist}(z,\mathcal{C}_{1})$.
C.2 Isotonic regression: mean preservation and majorization
Lemma 1 (Mean preservation; PAVA).
Let $\hat{f}\in\arg\min_{f\in\mathcal{M}_{\uparrow}}\sum_{i\in I}(y_{i}-f(s_{i}))^{2}$ be the isotonic fit (PAVA) on indices $I$. Then $\frac{1}{|I|}\sum_{i\in I}\hat{f}(s_{i})=\frac{1}{|I|}\sum_{i\in I}y_{i}$.
Lemma 2 (Dispersion reduction by majorization).
After sorting by $s$, the isotonic fitted vector $\hat{u}$ is a mean-preserving adjacent pooling of $y$; hence for any convex $\phi$, $\sum_{i}\phi(\hat{u}_{i})\leq\sum_{i}\phi(y_{i})$ (HardyLittlewoodPolya1952; MarshallOlkinArnold2011). In particular, with sample mean one, $\operatorname{Var}_{n}(\hat{u})\leq\operatorname{Var}_{n}(y)$ and $\mathrm{ESS}(\hat{u})\geq\mathrm{ESS}(y)$.
Proofs are standard; see Ayer1955PAVA; Barlow1972; RobertsonWrightDykstra1988; Banerjee2001.
C.3 Proof of Theorem 1 (surrogate EIF & variance drop)
Let $R^{\star}=\mathbb{E}[Y\mid S]$ and $m^{\star}(S)=\mathbb{E}[W_{\pi^{\prime}}\mid S]$. Under mean-sufficiency $\mathbb{E}[Y\mid X,A,S]=R^{\star}(S)$,
$V(\pi^{\prime})=\mathbb{E}\!\big[W_{\pi^{\prime}}R^{\star}\big]=\mathbb{E}\!\big[m^{\star}(S)R^{\star}(S)\big].$ (4)
Standard semiparametric calculations (projecting the unconstrained score onto the tangent space of the surrogate model) yield
$\phi_{\mathrm{sur}}(O;\pi^{\prime})=g^{\star}_{\pi^{\prime},R}(X)+m^{\star}(S)\big(R^{\star}-q^{\star}_{R}(X,A)\big)-V(\pi^{\prime}),$ (5)
with $q^{\star}_{R}(x,a)=\mathbb{E}[R^{\star}\mid x,a]$ and $g^{\star}_{\pi^{\prime},R}(x)=\mathbb{E}_{A\sim\pi^{\prime}(\cdot\mid x)}[q^{\star}_{R}(x,A)]$ (Bickel1993; VaartWellner2000). Since $m^{\star}$ is the $L^{2}$ projection of $W_{\pi^{\prime}}$ onto $L^{2}(\sigma(S))$, Pythagoras (or (3)) implies $\operatorname{Var}(\phi_{\mathrm{sur}})\leq\operatorname{Var}(\phi_{\mathrm{uncon}})$, strictly unless $W_{\pi^{\prime}}\in L^{2}(\sigma(S))$ and $R^{\star}$ is degenerate.
C.4 Proof of Theorem 2 (Efficiency via Model Restriction)
Let $\mathcal{M}$ be a baseline semiparametric model with tangent space $T(P)$ and canonical gradient $\phi^{\star}$. Let $\mathcal{M}_{\mathcal{C}}\subset\mathcal{M}$ be a restricted model with tangent space $T_{\mathcal{C}}(P)\subseteq T(P)$ and canonical gradient $\phi^{\star}_{\mathcal{C}}$.
The canonical gradient is defined as the projection of the pathwise derivative onto the tangent space. Since $T_{\mathcal{C}}(P)\subseteq T(P)$, we can decompose:
$\phi^{\star}=\phi^{\star}_{\mathcal{C}}+\phi^{\perp},$
where $\phi^{\perp}\in T(P)\cap T_{\mathcal{C}}(P)^{\perp}$ (the orthogonal complement of $T_{\mathcal{C}}(P)$ within $T(P)$).
By the Pythagorean theorem in Hilbert spaces:
$\|\phi^{\star}\|_{2}^{2}=\|\phi^{\star}_{\mathcal{C}}\|_{2}^{2}+\|\phi^{\perp}\|_{2}^{2}\geq\|\phi^{\star}_{\mathcal{C}}\|_{2}^{2},$
with equality iff $\phi^{\perp}=0$, i.e., $T_{\mathcal{C}}(P)=T(P)$ (no actual restriction).
Monotonicity for nested restrictions follows: if $T_{\mathcal{C}_{1}}(P)\supseteq T_{\mathcal{C}_{2}}(P)$ (i.e., $\mathcal{C}_{2}$ imposes stronger restrictions, yielding a smaller tangent space), then $\|\phi^{\star}_{\mathcal{C}_{2}}\|_{2}^{2}\leq\|\phi^{\star}_{\mathcal{C}_{1}}\|_{2}^{2}$.
For attainability, standard semiparametric arguments apply: one-step corrections or TMLE with cross-fitting yield regular estimators achieving the restricted efficiency bound (Bickel1993; VaartWellner2000; vanDerLaanRose2011TL).
C.5 Proof of Corollary 1
If $\sigma(S_{2})\subseteq\sigma(S_{1})$, then $L^{2}(\sigma(S_{2}))\subseteq L^{2}(\sigma(S_{1}))$. The surrogate model using $S_{1}$ has tangent space $T(S_{1})$, and the model using $S_{2}$ has $T(S_{2})\subseteq T(S_{1})$ when $\sigma(S_{2})\subseteq\sigma(S_{1})$. By Theorem 1 and the nested-tangent-space argument in Theorem 2, the canonical gradient in the smaller tangent space has variance no larger than in the larger one: $\operatorname{Var}(\phi_{\mathrm{sur}}(S_{1}))\leq\operatorname{Var}(\phi_{\mathrm{sur}}(S_{2}))$. Strictness fails only if the finer knowledge already holds, i.e., if $W_{\pi^{\prime}}$ is $\sigma(S_{2})$-measurable and $R^{\star}$ is degenerate.
C.6 Proof of 1 (Cal-IPS)
Write
$\widehat{V}_{\mathrm{IPS}}-V(\pi^{\prime})=(P_{n}-P)\!\big[m^{\star}(S)R^{\star}\big]+P\!\big[(\hat{W}_{\pi^{\prime}}-m^{\star})R^{\star}\big]+P\!\big[m^{\star}(\hat{f}(Z(S))-R^{\star})\big]+\mathrm{rem},$
where $\mathrm{rem}$ collects second-order sample-splitting terms. The empirical process term is $O_{p}(n^{-1/2})$; the second and third vanish by $L^{2}$-consistency of SIMCal-W and AutoCal-R (monotone or two-stage) and Cauchy–Schwarz; the remainder is $o_{p}(1)$ by cross-fitting. Finite-sample dispersion control follows from Lemma 2 and, if used, the blend cap $\rho\geq 1$; under the exact $\mathrm{IsoMeanOne}_{S}$ projection, $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})\geq\mathrm{ESS}(W_{\pi^{\prime}}^{\mathrm{m1}})$.
C.7 Proof of Theorem 3 (DR-CPO $\sqrt{n}$ limits)
With cross-fitted nuisances and $R^{\mathrm{OOF}}$ along the IF path,
$\widehat{V}_{\mathrm{DR}}-V(\pi^{\prime})=(P_{n}-P)\!\big[\phi_{\mathrm{sur}}\big]+P\!\big[(\hat{W}_{\pi^{\prime}}-m^{\star})\{q^{\star}_{R}-\hat{q}\}\big]+o_{p}(n^{-1/2}).$
The second term is $o_{p}(n^{-1/2})$ by the one-of-two product-rate condition $\|\hat{W}_{\pi^{\prime}}-m^{\star}\|_{2}\cdot\|\hat{q}-q^{\star}_{R}\|_{2}=o_{p}(n^{-1/2})$ and cross-fitting; the central limit theorem yields the limit variance $\operatorname{Var}(\phi_{\mathrm{sur}})$.
C.8 Proof of Theorem 4 (budgeted bound)
Let $\mathcal{W}_{\rho}=\{m:\mathbb{E}[m]=1,\ m\uparrow S,\ \mathbb{E}[(m-1)^{2}]\leq\rho\,\mathbb{E}[(m^{\star}-1)^{2}]\}$. Intersecting the surrogate tangent space with the linear span induced by $m\in\mathcal{W}_{\rho}$ replaces $m^{\star}$ by its $L^{2}(P_{S})$ projection $m^{\star}_{\rho}=\Pi_{\mathcal{W}_{\rho}}(m^{\star})$ in $\phi_{\mathrm{sur}}$, giving $\phi^{(\rho)}$. Monotonicity in $\rho$ follows from nested convex sets $\mathcal{W}_{\rho_{1}}\subseteq\mathcal{W}_{\rho_{2}}$ for $\rho_{1}\leq\rho_{2}$, and $\lim_{\rho\to\infty}\phi^{(\rho)}=\phi_{\mathrm{sur}}$. If SIMCal-W converges to $m^{\star}_{\rho}$ (with the same cap), DR-CPO attains $\operatorname{Var}(\phi^{(\rho)})$ by the same one-of-two rate argument.
C.9 Proof of Theorem 5 (IF-space stacking) and Corollary 2
Let $\{\widehat{V}^{(e)}\}$ be regular and asymptotically linear with centered IFs $\{\phi^{(e)}\}$. Set $\Phi=[\phi^{(e)}]$ and $\hat{\Sigma}=(1/n)\Phi^{\top}\Phi+\lambda I$. A uniform law of large numbers on the simplex $\Delta$ yields $\hat{\Sigma}\to\Sigma$ uniformly; by argmin continuity, $\hat{\alpha}\to\alpha^{\star}\in\arg\min_{\alpha\in\Delta}\alpha^{\top}\Sigma\,\alpha$. Hence $\widehat{V}^{(\hat{\alpha})}$ is asymptotically linear with IF $\phi^{(\alpha^{\star})}=\sum_{e}\alpha^{\star}_{e}\phi^{(e)}$ and variance $\min_{\alpha\in\Delta}\alpha^{\top}\Sigma\,\alpha\leq\min_{e}\Sigma_{ee}$. For Corollary 2: if $\mathrm{rank}(\Phi)=r$, then the feasible IF combinations lie in an $r$-dimensional affine subspace; by Carathéodory’s theorem, any point in $\mathrm{conv}\{\phi^{(e)}\}$ admits a representation using at most r+1r{+}1 extreme points.
C.10 Proof of 2 (OUA jackknife)
[…]
C.11 Auxiliary lemmas used in the main proofs
Lemma 3 (OOF mean preservation for AutoCal-R).
Let $K$ be fixed and let $R^{\mathrm{OOF}}_{i}=\hat{f}^{(-F(i))}(Z(S_{i}))$ be OOF predictions from AutoCal-R (either mode). Then $P_{n}[R^{\mathrm{OOF}}]-P_{n}[Y]=o_{p}(1)$ and $P[R^{\mathrm{OOF}}-R^{\star}]=o_{p}(1)$ under $L^{2}(P_{S})$-consistency of $\hat{f}$.
Lemma 4 (Second-order remainder for DR with cross-fitting).
Let $\widehat{V}_{\mathrm{DR}}$ be DR-CPO with cross-fitted $(\hat{q},\hat{g}_{\pi^{\prime}})$ and calibrated $\hat{W}_{\pi^{\prime}}$. Then
$\widehat{V}_{\mathrm{DR}}-V(\pi^{\prime})-(P_{n}-P)\phi_{\mathrm{sur}}=P\!\big[(\hat{W}_{\pi^{\prime}}-m^{\star})(q^{\star}_{R}-\hat{q})\big]+o_{p}(n^{-1/2}),$
and the bracketed term is $o_{p}(n^{-1/2})$ under $\|\hat{W}_{\pi^{\prime}}-m^{\star}\|_{2}\cdot\|\hat{q}-q^{\star}_{R}\|_{2}=o_{p}(n^{-1/2})$.
Lemma 5 (Guard stability).
Let $W^{\mathrm{stack}}$ be the OOF-stacked candidate and $W^{\mathrm{m1}}_{\pi^{\prime}}$ the mean-one baseline. For $\rho\geq 1$, define $\alpha=\min\!\big\{1,\sqrt{\rho\,\operatorname{Var}(W^{\mathrm{m1}}_{\pi^{\prime}})/\operatorname{Var}(W^{\mathrm{stack}})}\big\}$ and $W^{\mathrm{blend}}=1+\alpha(W^{\mathrm{stack}}-1)$. Then $\operatorname{Var}(W^{\mathrm{blend}})=\alpha^{2}\operatorname{Var}(W^{\mathrm{stack}})\leq\rho\,\operatorname{Var}(W^{\mathrm{m1}}_{\pi^{\prime}})$, and the subsequent mean-one isotonic projection cannot increase empirical variance by Lemma 2.
Implementation note. The reference implementation uses a simpler absolute-cap formula $\alpha=\min\{1,\sqrt{\rho/\operatorname{Var}(W^{\mathrm{stack}})}\}$ and omits the final isotonic re-projection (normalizing to mean one instead). This relaxation empirically yields similar ESS improvements but does not carry the deterministic guarantee of Lemma 5.
Lemma 6 (Firm non-expansiveness of projections).
Metric projections onto closed convex sets in Hilbert spaces are firmly non-expansive: $\|\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y)\|^{2}\leq\langle\Pi_{\mathcal{C}}(x)-\Pi_{\mathcal{C}}(y),\,x-y\rangle$. Hence compositions of DbP modules (reward, weight, IF-space projections) are non-expansive. See, e.g., BauschkeCombettes2017.
Remarks on dependence.
If logs exhibit serial or cluster dependence, the IF CLTs can be replaced by block/stationary bootstrap arguments; our reported intervals can include a dependence-robust alternative (App. A.8).
C.12 Coverage-Limited Efficiency (CLE) Diagnostic
We derive the CLE diagnostic stated in Equation 1. Let $\mathcal{T}\subset\mathcal{X}\times\mathcal{A}$ be a target-relevant region with $\alpha=P_{\pi^{\prime}}(\mathcal{T})$ and $\beta=P_{\pi_{0}}(\mathcal{T})$.
Lemma 7 (Weight second moment on $\mathcal{T}$).
For the normalized conditional distributions $\pi^{\prime}_{\mathcal{T}}$ and $\pi_{0,\mathcal{T}}$,
$\mathbb{E}_{\pi_{0}}\!\big[W_{\pi^{\prime}}^{2}\,\mathbf{1}_{\mathcal{T}}\big]=\frac{\alpha^{2}}{\beta}\bigl(1+\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})\bigr).$
Proof.
By definition of $\chi^{2}$ divergence for the normalized conditionals: $\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})=\int_{\mathcal{T}}\frac{(d\pi^{\prime}_{\mathcal{T}})^{2}}{d\pi_{0,\mathcal{T}}}-1=\frac{\beta}{\alpha^{2}}\int_{\mathcal{T}}\frac{(d\pi^{\prime})^{2}}{d\pi_{0}}-1=\frac{\beta}{\alpha^{2}}\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}\mathbf{1}_{\mathcal{T}}]-1$. Rearranging gives the stated identity. ∎
Proposition 4 (CLE diagnostic for IPS-style estimators).
For any IPS-style estimator $\hat{\Psi}_{\mathrm{IPS}}$ of $V(\pi^{\prime})=\mathbb{E}_{\pi^{\prime}}[Y]$ based on importance-weighted outcomes,
$\operatorname{SE}(\hat{\Psi}_{\mathrm{IPS}})\;\geq\;\frac{\sigma_{\mathcal{T}}\,\alpha}{\sqrt{\beta\,n}}\,\sqrt{1+\chi^{2}(\pi^{\prime}_{\mathcal{T}}\|\pi_{0,\mathcal{T}})},$
where $\sigma_{\mathcal{T}}^{2}:=\mathrm{ess\,inf}_{(x,a)\in\mathcal{T}}\operatorname{Var}(Y\mid X{=}x,A{=}a)$ is the minimal outcome noise in $\mathcal{T}$.
Proof.
We lower bound $\operatorname{Var}_{\pi_{0}}(W_{\pi^{\prime}}Y)$ using the Law of Total Variance, conditioning on $(X,A)$:
$\operatorname{Var}_{\pi_{0}}(W_{\pi^{\prime}}Y)=\mathbb{E}_{\pi_{0}}[\operatorname{Var}(W_{\pi^{\prime}}Y\mid X,A)]+\operatorname{Var}_{\pi_{0}}(\mathbb{E}[W_{\pi^{\prime}}Y\mid X,A]).$
Since $W_{\pi^{\prime}}$ is $(X,A)$-measurable:
$\operatorname{Var}(W_{\pi^{\prime}}Y\mid X,A)$ $=W_{\pi^{\prime}}^{2}\operatorname{Var}(Y\mid X,A),$
$\mathbb{E}[W_{\pi^{\prime}}Y\mid X,A]$ $=W_{\pi^{\prime}}\mathbb{E}[Y\mid X,A].$
Dropping the non-negative second term:
$\operatorname{Var}_{\pi_{0}}(W_{\pi^{\prime}}Y)\geq\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}\operatorname{Var}(Y\mid X,A)].$
Restricting to $\mathcal{T}$ (non-negative integrand) and applying $\operatorname{Var}(Y\mid X,A)\geq\sigma_{\mathcal{T}}^{2}$ on $\mathcal{T}$:
$\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}\operatorname{Var}(Y\mid X,A)]\geq\sigma_{\mathcal{T}}^{2}\mathbb{E}_{\pi_{0}}[W_{\pi^{\prime}}^{2}\mathbf{1}_{\mathcal{T}}].$
Applying Lemma 7:
$\operatorname{Var}_{\pi_{0}}(W_{\pi^{\prime}}Y)\geq\frac{\sigma_{\mathcal{T}}^{2}\alpha^{2}}{\beta}(1+\chi^{2}).$
The asymptotic variance of IPS is $\operatorname{Var}_{\pi_{0}}(W_{\pi^{\prime}}Y)/n$; taking square roots yields the stated bound. ∎
Interpretation.
The CLE diagnostic has three multiplicative factors: (i) the coverage penalty $\alpha/\sqrt{\beta}$, which explodes when the logger rarely visits target-typical regions ($\beta\ll\alpha$); (ii) the shape mismatch $\sqrt{1+\chi^{2}}$, which inflates the floor even with good coverage if the target concentrates differently within $\mathcal{T}$; (iii) the standard noise term $\sigma_{\mathcal{T}}/\sqrt{n}$.
Corollary (ESS form).
Using $\mathrm{ESS}(W_{\pi^{\prime}})=n/(1+\operatorname{CV}^{2}(W_{\pi^{\prime}}))$ and the relationship between $\chi^{2}$ and ESS on $\mathcal{T}$:
$\operatorname{SE}(\hat{\Psi})\;\geq\;\frac{\sigma_{\mathcal{T}}}{\sqrt{\mathrm{ESS}_{\mathcal{T}}}},$
where $\mathrm{ESS}_{\mathcal{T}}$ is the effective sample size restricted to $\mathcal{T}$. This shows that high global ESS is insufficient: what matters is ESS in target-relevant regions.
C.13 Assumptions ledger
For any CJE deployment, record and validate the following assumptions:
A0 (Bridge Assumption).
$\mathbb{E}[Y^{*}\mid X,A]=\mathbb{E}[Y\mid X,A]$. The operational oracle $Y$ is sufficiently aligned with the idealized deliberation oracle $Y^{*}$ such that optimizing for $Y$ approximates optimizing for $Y^{*}$.
Validation: Bridge Validation Protocol (BVP) with predictive transportability experiments, construct validity audits, and stability monitoring.
S1 (Prentice Sufficiency).
$\mathbb{E}[Y\mid X,A,S]=f(S,X)$ a.s. on $\operatorname{supp}(\pi_{0}\cup\Pi_{\mathrm{eval}})$. The surrogate $S$ captures all relevant information about the oracle $Y$ given covariates $X$.
Testable implication: Residuals $Y-\hat{f}(S,X)$ should be mean-zero and uncorrelated with $A$ conditional on $(X,S)$.
S2 (Transportability).
The calibration function $f$ transports across environments $g\in\mathcal{G}$: $\mathbb{E}_{g}[Y\mid X,A,S]=f(S,X)$ for all $g$ (PearlBareinboim2014).
Graphical test: S-admissibility requires that no selection node points into $Y$ given $S$ in the selection diagram with incoming arrows to $A$ removed.
Testable implication: Calibration curves should be stable across (i) policy subgroups, (ii) time periods, and (iii) covariate strata.
S3 (Overlap/Positivity).
$\pi^{\prime}(a\mid x)>0\Rightarrow\pi_{0}(a\mid x)>0$ a.s. Required for off-policy estimation.
Diagnostic: Target-Typicality Coverage (TTC). If TTC <0.7<0.7, expect logs-only IPS to fail.
L1 (Oracle MAR).
$L\perp Y\mid(X,A,S)$, where $L\in\{0,1\}$ indicates oracle labeling. Oracle labeling is ignorable conditional on observed surrogates.
Testable implication: Oracle-labeled and unlabeled samples should have similar $(X,A,S)$ distributions.
L2 (Oracle Positivity).
$P(L=1\mid X,A,S)>0$ on the support where calibration will be applied.
Diagnostic: Check for gaps in oracle labeling across the $(X,A,S)$ space.
C.14 What the bounds do not include
The surrogate and budgeted information bounds describe model limits: they do not include (i) finite-sample dispersion control from the sample cap (we encode it population-wise via $\rho$) or (ii) the oracle first-stage uncertainty (added separately by OUA).
C.15 Citations for technical ingredients
Semiparametric efficiency and one-step/TMLE: Bickel1993; VaartWellner2000; vanDerLaanRose2011TL; Tsiatis2006.
Isotonic regression / PAVA and order-restricted inference: Ayer1955PAVA; Barlow1972; RobertsonWrightDykstra1988; Banerjee2001.
Majorization theory: HardyLittlewoodPolya1952; MarshallOlkinArnold2011.
Calibration for OPE: KallusMao2020Surrogates; vanDerLaanLinCaroneLuedtke2025StabIPW; vanDerLaanLuedtkeCarone2024DRCalibration.
Jackknife/bootstraps for dependence: Kunsch1989BlockBootstrap; PolitisRomano1994StationaryBootstrap.
Projections and non-expansive maps: BauschkeCombettes2017.
Appendix D Diagnostics, Gates, and Reporting (details)
This appendix formalizes the diagnostics used in CJE, the associated ship/stop gates, and the reporting ledger. The main text shows a compact panel per policy; here we provide precise formulas, defaults, and recommended thresholds. Unless stated otherwise, expectations and variances are empirical over the evaluation cohort (global SNIPS normalization).
D.1 Weight behavior & overlap
Effective sample size (ESS).
For nonnegative weights,
$\mathrm{ESS}(W)=\frac{\big(\sum_{i}W_{i}\big)^{2}}{\sum_{i}W_{i}^{2}}.$
Under global mean-one normalization (SNIPS), $\sum_{i}W_{i}=n$, so
$\frac{\mathrm{ESS}(W)}{n}=\frac{1}{1+\mathrm{CV}^{2}(W)},\qquad\mathrm{CV}^{2}(W)=\operatorname{Var}(W)\ \text{when }\mathbb{E}[W]=1.$
Report the ESS fraction $\mathrm{ESS}(W)/n$ and the multiplicative uplift $\mathrm{ESS}(\hat{W}_{\pi^{\prime}})/\mathrm{ESS}(W^{\mathrm{m1}}_{\pi^{\prime}})$.
Max-weight share.
$\max_{i}W_{i}\big/\sum_{j}W_{j}$ flags single-row dominance; display alongside the empirical 99.5th percentile of $W$.
Tail index (Hill) and CCDF.
For the top-$k$ order statistics $W_{(1)}\geq\cdots\geq W_{(k)}$,
$\hat{\alpha}^{-1}(k)=\frac{1}{k}\sum_{j=1}^{k}\log\!\left(\frac{W_{(j)}}{W_{(k)}}\right),\qquad k\in\mathcal{K},$
swept over a stability grid $\mathcal{K}$ (e.g., 1–5% of $n$). Plot $\hat{\alpha}(k)$ with a band over the plateau region (median and IQR over $\mathcal{K}$), and the empirical CCDF of $W$ on a log–log scale.
Overlap in judge space.
Let $p_{S|\pi_{0}}$ and $p_{S|\pi^{\prime}}$ denote the (binned) densities of $S$ under $\pi_{0}$ and $\pi^{\prime}$:
$A_{B}=\sum_{b}\sqrt{p_{b}(\pi_{0})\,p_{b}(\pi^{\prime})},\qquad D_{B}=-\log A_{B}.$
Overlay an $S$-binned heatmap of $\log W$ to localize regions of poor overlap.
D.2 Judge calibration, coverage, and drift
Reliability diagram.
Partition $S$ into $B$ bins; for bin $b$, plot the bin mean of $R=\hat{f}(Z(S))$ against the oracle mean of $Y$, with 95% binomial intervals. Report a Brier-style reliability term and the OOF RMSE, with a 1-SE model-selection overlay (monotone vs. two-stage).
Coverage badge.
Estimate the fraction of evaluation mass outside the oracle $S$ range:
$\texttt{OutOfRange}=\widehat{\Pr}_{\pi^{\prime}}\!\big(S<S_{\min}^{\text{orc}}\ \text{or}\ S>S_{\max}^{\text{orc}}\big).$
Also report boundary flatness (slope of $\hat{f}$ in the lowest/highest oracle decile). Large OutOfRange together with flat boundaries triggers Limited Calibration Support and the REFUSE-LEVEL gate.
Rank drift (optional anchor).
Given a fixed anchor set of $(X,A)$ pairs scored over time, compute Kendall’s $\tau$ between historical and current judge rankings with a permutation $p$-value. Change detection on residuals can be monitored via CUSUM/EWMA with FDR control across anchors.
D.3 DR orthogonality and decomposition
Orthogonality score.
Let $U_{i}=\hat{W}_{\pi^{\prime},i}\big(R^{\mathrm{OOF}}_{i}-\hat{q}^{\mathrm{OOF}}(X_{i},A_{i})\big)$ and $\bar{U}=n^{-1}\sum_{i}U_{i}$. Form a Wald CI for $\bar{U}$ using the standard error $\sqrt{\widehat{\operatorname{Var}}(U)/n}$ (or a cluster-/block-robust analogue). Report $\bar{U}$ and its CI; near-zero indicates successful orthogonality.
DM–IPS decomposition.
Display $\widehat{V}_{\mathrm{DM}}=n^{-1}\sum_{i}\hat{g}_{\pi^{\prime}}(X_{i})$ and $\widehat{V}_{\mathrm{Aug}}=n^{-1}\sum_{i}U_{i}$, with CIs and the empirical correlation between their per-row contributions.
D.4 Uncertainty: IF variance and OUA addition
For any estimator with centered IF contributions $\{\phi_{i}\}_{i=1}^{n}$,
$\widehat{\operatorname{Var}}_{\mathrm{main}}=\frac{1}{n}\,\widehat{\operatorname{Var}}(\phi_{i}),\qquad\widehat{\operatorname{Var}}_{\mathrm{total}}=\widehat{\operatorname{Var}}_{\mathrm{main}}+\widehat{\operatorname{Var}}_{\mathrm{cal}},$
with $\widehat{\operatorname{Var}}_{\mathrm{cal}}$ from the oracle jackknife (App. A.7). Report the OUA share $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ and, optionally, a dependence-robust alternative (below).
Dependence-robust SEs.
When time/cluster dependence is suspected, also report: (i) cluster-robust sandwich SEs when a cluster id (e.g., session/user) is available; and (ii) block/stationary bootstrap intervals (block length chosen by a simple variance-stability sweep) (Kunsch1989BlockBootstrap; PolitisRomano1994StationaryBootstrap).
D.5 Multiplicity for many-policy comparisons
For contrasts $\Delta_{p}=\widehat{V}(\pi^{\prime}_{p})-\widehat{V}(\pi^{\star})$, compute Wald $p$-values and apply BH at level $q\in[0.05,0.2]$; BY can be used under strong dependence. Provide a pairwise win matrix with FDR marks and Kendall’s $\tau$ over policy means (all policies).
D.6 Policy-wise mean transport diagnostic
For each target policy $\pi^{\prime}$ (or more generally, each evaluation environment such as a time period or subgroup), we test whether the calibration function $f(S,X)$ learned on the base policy remains mean-unbiased.
Test procedure.
Given an oracle slice under policy $\pi^{\prime}$, compute the residual $\varepsilon_{\pi^{\prime}}=Y-f(S,X)$ and test
$H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[\varepsilon_{\pi^{\prime}}]=0\quad\text{vs}\quad H_{1,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[\varepsilon_{\pi^{\prime}}]\neq 0$
using a one-sample $t$-test. When testing multiple policies, apply Bonferroni or BH correction across environments.
Interpretation (Proposition 3).
By 3, $H_{0,\pi^{\prime}}$ holds if and only if the surrogate-estimated policy value equals the oracle value: $\mathbb{E}_{\pi^{\prime}}[f(S,X)]=\mathbb{E}_{\pi^{\prime}}[Y]$. Rejecting $H_{0,\pi^{\prime}}$ implies that surrogate-only evaluation is systematically biased for policy $\pi^{\prime}$: the calibration does not transport.
OUA-corrected inference.
To account for uncertainty in estimating $f$, the confidence interval for the mean residual should incorporate OUA variance (Section 3.5). In practice, we refit $f^{(-k)}$ omitting each oracle fold and compute residuals under each refit, propagating calibration uncertainty into the test statistic.
Actions on failure.
If a policy fails the mean transport test: (i) flag surrogate-only value estimates for that policy as biased; (ii) report the estimated bias ($\hat{\Delta}=\bar{\varepsilon}_{\pi^{\prime}}$) and its direction; (iii) either recalibrate using policy-specific oracle data or fall back to oracle-only evaluation for that cell; (iv) rankings across policies that all pass may still be valid even if absolute levels are not.
D.7 Gates: thresholds and actions
Table 6: Default gates (suggested; tighten for high-stakes launches).
Gate Default Action if failed
Overlap $\mathrm{ESS}/n\geq 0.30$ (${\leq}3.3{\times}$ var); Hill $\hat{\alpha}\geq 2$ (finite var); $A_{B}\geq 0.85$ ($H{\leq}0.39$); TTC $\geq 0.70$ Use overlap weights or cohort restriction; if TTC<0.70{}<0.70, prefer Direct over IPS; report with warning if $\hat{\alpha}\in[1,2)$; do not ship offline conclusions if $\hat{\alpha}<1$
Judge Reliability band covers diagonal at knots; no persistent drift alarms Refresh/extend oracle slice; switch to two-stage index; re-validate
Identification OutOfRange $\leq\eta$ (default $\eta{=}5\%$; worst-case bias ${\leq}\eta$) or non-flat boundaries Flag Limited Calibration Support; set REFUSE-LEVEL: report rankings + partial-ID only
DR Orthogonality CI includes 0; no NaNs; residual tails acceptable Strengthen nuisances/cross-fitting; fall back to stabilized IPS as a diagnostic
Multiplicity FDR control applied when $|\Pi^{\prime}|>5$ Report adjusted $p$-values; avoid uncorrected winner claims
Cap Guard rarely engaged; CI width not sensitive to $\rho$ If guard active on $50\%$ folds or sensitivity high, show cap curve and prefer overlap weights/restriction
Threshold calibration.
[…]
REFUSE-LEVEL procedure.
When Identification fails: (i) gray-out level estimates; (ii) highlight OutOfRange and boundary flatness; (iii) report rank-only conclusions with conservative relative CIs; (iv) recommend targeted labeling in uncovered $S$ regions.
Table 7: Assumptions Ledger: Quick Reference. For each assumption, we list the estimators it applies to, how to test it, and recommended mitigations if violated.
Assumption Applies To How to Test If Violated
[…]
Mean sufficiency All Reliability diagram; regional residuals centered at zero Two-stage fallback with covariates; refresh oracle slice
Monotonicity in $S$ AutoCal-R OOF model selection: monotone vs. two-stage; 1-SE rule Use two-stage with response length covariate
Calibration coverage All OutOfRange $\leq 5\%$ (worst-case bias ${\leq}5\%$); non-flat boundaries REFUSE-LEVEL; targeted labeling in uncovered regions
Shared prompts Direct Paired design check (same prompt set across policies) Cannot use Direct; must use OPE methods
Transportability Direct Residual test: $\mathbb{E}[Y-f(S,X)]=0$ per policy (Bonferroni) Recalibrate per policy; collect policy-specific oracle data
D.8 Planner: MDE and label/log budgets
Given two independent estimates with equal SE $\widehat{\mathrm{SE}}$, the two-sided 95% test at 80% power has
$\mathrm{MDE}_{80\%}=(z_{0.8}+z_{0.975})\sqrt{2}\;\widehat{\mathrm{SE}}$.
We tabulate $\widehat{\mathrm{SE}}$ versus (n,m/n)(n,\,m/n) (labels per log) using Stacked-DR with OUA and annotate “iso-cost” lines for the label budget.
D.9 Reporting ledger (per policy/cohort)
Persist: (i) calibrator mode, OOF risk by tertiles, knots/levels (hash); (ii) SIMCal–W maps, stacking weights $\hat{\beta}$, guard $\rho$ and blend $\alpha$; (iii) ESS fraction, max-weight share, Hill index band, $A_{B}$; (iv) DR orthogonality score and CI; DM–IPS split; (v) OUA trace $\{\widehat{V}^{(-k)}\}$ and variance breakdown; (vi) filter counts (e.g., TF gaps) and an inclusion manifest of $x$-ids; (vii) multiplicity control (family, $q$, adjusted $p$).
D.10 Visualization primitives (for reproducible panels)
  • ESS/tails strip: bars for ESS fraction (baseline vs. SIMCal–W); dot for max-weight share; Hill band.
  • $S$-overlap heatmap: density of $S$ under $\pi_{0}$ vs. $\pi^{\prime}$ with overlaid $\log W$; annotate $A_{B}$.
  • Reliability panel: bin means of $(R,Y)$ with 95% CIs; mode card (monotone vs. two-stage; OOF RMSE).
  • Orthogonality panel: point/CI for $\bar{U}$; DM–IPS bars with CIs and correlation.
  • Uncertainty ring: pie of $\widehat{\operatorname{Var}}_{\mathrm{cal}}/\widehat{\operatorname{Var}}_{\mathrm{total}}$ (OUA share).
D.11 Optional: dependence–robust implementation details
Cluster-robust SEs: if a cluster id $c(i)$ is available, $\widehat{\operatorname{Var}}_{\mathrm{CR}}=n^{-2}\sum_{c}\big(\sum_{i\in c}\phi_{i}\big)\big(\sum_{i\in c}\phi_{i}\big)^{\!\top}$, with finite-sample correction. Stationary bootstrap: sample blocks of geometric length $\ell\sim\mathrm{Geom}(p)$ glued to length $n$; form the bootstrap distribution of $\hat{\psi}$ (or of $n^{-1/2}\sum\phi_{i}$) and report percentile or $t$-based bands.
D.12 Compact gate pseudo-logic
Algorithm 7 Gate logic (per policy)
1: Compute weight/tail metrics: ESS fraction, max-share, Hill band; compute $A_{B}$; judge reliability/coverage; orthogonality score; OUA share
2:if $\mathrm{ESS}/n<0.30$ or median Hill<2<2 or $A_{B}<0.85$ then
3:  Flag Overlap (warn; restrict or use overlap weights)
4:end if
[…]
6:  REFUSE-LEVEL $\leftarrow$ TRUE
7:end if
8:if Orthogonality CI excludes 0 then
9:  Flag DR; strengthen nuisances/cross-fitting
10:end if
11:if Cap engaged on $50\%$ folds or CI sensitivity to $\rho$ high then
12:  Show cap–sensitivity; prefer overlap weights/restriction
13:end if
14: Apply multiplicity control (BH/BY) when $|\Pi^{\prime}|>5$
Appendix E Implementation, Engineering, and Reproducibility
This appendix enumerates the concrete artifacts needed to reproduce CJE end–to–end: a minimal logging schema, a teacher–forcing (TF) contract with conformance checks, fold construction, numerics, persisted outputs, and a lightweight resource model. No packages beyond the ICLR style file are required.
E.1 Minimal logging schema (storage–agnostic)
Each row corresponds to one prompt–continuation pair under the fixed logger $\pi_{0}$. We persist only what is necessary to reconstruct SNIPS/IPS weights and judge scores.
Table 8: Columns required for CJE. Columnar formats (Parquet) are convenient but not required.
Field Type Description
x_id string Stable identifier (hash of normalized prompt ++ cohort)
prompt bytes/string Canonicalized $X$ (tokenizer ++ normalization recorded)
continuation bytes/string Realized $A$ under $\pi_{0}$ (full sequence)
tokens int[] Token ids for $A$ under each model’s TF tokenizer
logp_pi0 float[] Per–token $\log p_{\pi_{0}}(a_{t}\!\mid\!h_{t})$ across AA
judge_S float/json Scalar judge score $S=s(X,A)$ (or struct of sub–scores)
judge_cfg json Judge rubric, decoding params, model snapshot hash
run_cfg json $\pi_{0}$ engine tag, decoding params, checkpoint hash, seed
fold_id int Deterministic fold (F(x_id)F(x\_id); see §E.3)
cohort string Optional slice label (time window, traffic source, etc.)
TF cache (per target $\pi^{\prime}$).
[…]
$W^{\mathrm{m1}}_{\pi^{\prime}}\;=\;\exp\!\Big(\texttt{logW}-\mathrm{logsumexp}(\texttt{logW})+\log n\Big),$
i.e., a single global denominator that enforces sample–mean–one. Rows with missing/invalid TF are filtered and recorded in a ledger.
E.2 Teacher forcing: contract and conformance
We require a single–call, chat–native TF API that returns per–token and summed $\log p_{\pi}(A\!\mid\!X)$ under a fixed template, tokenizer, and snapshot. Client–side checks:
[…]
  • Additivity. Also return $\log p_{\pi}(X)$ and $\log p_{\pi}(X{+}A)$ and verify $\log p_{\pi}(X{+}A)\approx\log p_{\pi}(X)+\log p_{\pi}(A\!\mid\!X)\leq\log p_{\pi}(X)$. Violation $\Rightarrow$ discard row (and log it).
  • Template/tokenizer provenance. Return immutable hashes; reject moving aliases.
  • Masking. If safety masks are applied, return mask bits and a renormalized flag; prefer an evaluation–only path without hidden renormalization.
Conformance snippet (pseudo).
lp = TF(model_id, template_id, X, A)
assert same_bits(lp.sum, sum(lp.per_token), tol)
lpX, lpXA = TF_logp(X), TF_logp(X+A)
assert abs(lpXA - (lpX + lp.sum)) < eps and lp.sum <= 0
E.3 Folds and cross–fitting
We use K=5K{=}5 folds by default. The fold map $F(i)$ is a stable hash of x_id modulo $K$, ensuring that all modules (AutoCal–R, SIMCal–W, DR nuisances) share identical OOF boundaries. Oracle folds are derived by intersecting $F(i)$ with $L_{i}{=}1$. The hash rule is $F(x\_id){=}\mathrm{hash}(x\_id)\bmod 5$.
E.4 Numerics and stability
  • Ratios in log–space. Keep $\log W$ until forming the global mean–one normalization; use a single logsumexp for the denominator.
  • Center residuals. For stacking objectives and covariances, center $\Delta$ and drop NaNs/inf at ingestion.
  • Variance estimates. Use Welford’s online formulas for high dynamic range; add a tiny ridge ($\lambda\!\in\![10^{-10},10^{-6}]$) to covariance matrices.
  • PAVA. Run once per fold after sorting by $S$. The reference implementation enforces mean-one via multiplicative rescaling ($W/\bar{W}$), which preserves nonnegativity but may slightly distort isotonicity at extremes. Note: The exact constrained projection $\mathrm{IsoMeanOne}_{S}$ defined in Section A.3 preserves both monotonicity and mean-one simultaneously; the deterministic majorization guarantee strictly applies only to that exact operator. We empirically observe similar ESS improvements with multiplicative rescaling.
  • Guard (absolute variance cap). Default $\rho{=}1$; compute $\alpha=\min\!\big\{1,\,\sqrt{\rho/\operatorname{Var}(W^{\mathrm{stack}})}\big\}$; blend and normalize to mean–one. Persist whether the guard engaged. Note: This is an absolute cap ($\operatorname{Var}\leq\rho$); the theoretical Lemma 5 uses a relative bound ($\operatorname{Var}\leq\rho\cdot\operatorname{Var}(W^{\mathrm{m1}})$).
Reference code note. A minimal reference implementation uses global isotonic fits and computes the stacking covariance on those same in-sample fits (no OOF); production code should use cross-fitting as described above.
E.5 Persisted artifacts (per policy/cohort)
  • Calibrator: mode (monotone vs. two–stage), OOF RMSE (overall ++ tertiles), knots/levels (hash), OOF vs. pooled predictions.
  • Weights: isotonic merge metadata, $S$ orientation (up/down), stacking weights $\hat{\beta}$ including the identity/baseline candidate, guard $\rho$ and blend $\alpha$, final mean–one check.
  • Estimators: point estimates, centered IF vectors’ hashes, $\widehat{\operatorname{Var}}_{\mathrm{main}}$, orthogonality score and CI, dependence–robust SEs (if used).
  • OUA: $\{\widehat{V}^{(-k)}\}_{k=1}^{K}$, $\widehat{\operatorname{Var}}_{\mathrm{cal}}$, $\widehat{\operatorname{Var}}_{\mathrm{total}}$.
  • Diagnostics: ESS fraction, max–weight share, Hill band, SS–overlap ($A_{B}$), coverage badge, gate statuses.
  • Ledger: counts by filter reason (TF gaps, moderation, timeouts), x_id inclusion manifest.
E.6 Reference run order (pseudocode)
# 0) Build TF cache for each pi’ (one pass per policy)
build_tf_cache --policies <list> --dataset logs.parquet --out tf_cache.parquet
# 1) Reward calibration (cross-fitted; auto monotone vs two-stage)
autocal_r --oracle oracle.parquet --folds 5 --out rewards.parquet
# 2) SIMCal-W per policy (OOF project->stack->cap->translate-to-mean-one)
simcal_w --tf-cache tf_cache.parquet --scores S.parquet --rho 1.0 --folds 5 \
         --out weights.parquet
# 3) Estimation + IFs (Cal-IPS / DR-CPO)
estimate --rewards rewards.parquet --weights weights.parquet --folds 5 \
         --out estimates.parquet
# 4) IF-Stack (optional)
stack --estimates estimates.parquet --out stacked.parquet
# 5) OUA jackknife
oua --oracle-folds 5 --pipeline-config cfg.yaml --out variance.parquet
# 6) Report (diagnostics, gates, CIs)
report --inputs *.parquet --figs figs/ --out report.html
E.7 Compute and resource model
Let $n$ be prompts, $\bar{T}$ the mean continuation length, and $|\Pi^{\prime}|$ the number of candidate policies.
  • TF cache. $O(|\Pi^{\prime}|\,n\,\bar{T})$ forward tokens; microbatch by length; near–linear scaling across GPUs.
  • SIMCal–W. $O(n\log n)$ for sort ++ $O(n)$ for PAVA per fold; covariance/stacking are tiny ($3{\times}3$).
  • DR-CPO. If $\hat{g}_{\pi^{\prime}}$ uses one rollout per $(X,\pi^{\prime})$, add $O(|\Pi^{\prime}|\,n\,\bar{T}^{\prime})$ tokens once; a light smoother amortizes Monte Carlo noise.
  • OUA. $K$ refits of AutoCal–R and re–runs of the pipeline; cache features to avoid recomputation.
E.8 Determinism, versioning, and privacy
Determinism: fix seeds at engine, dataloader, and sampler; record random states in run_cfg; serialize fold maps (internal).
Versioning: record immutable hashes for model weights, tokenizer, and template; pin checkpoints.
Privacy: encrypt prompts/continuations at rest; hash x_id with salt; public artifacts include only aggregates/diagnostics and redacted IDs.
E.9 Exact prompts used in experiments
Policy system prompts.
All policies use Llama 3.3 70B except premium (Llama 3.1 405B), served via Fireworks AI (https://fireworks.ai). Temperature = 0.7 for all; no top-pp/top-$k$ truncation, so TF logprobs represent the exact sampling distribution.
  • base, clone, premium: “You are a helpful assistant.”
  • parallel_universe_prompt: “Imagine parallel universes where you vary your responses and can observe which one improves the user’s life the most. Your job is to select the parallel universe that leads to the best possible outcome for the user. Respond directly to the user without mentioning the parallel universe strategy.”
  • unhelpful: “You are an unhelpful assistant that deliberately confuses and misleads the user.”
Judge and oracle evaluation prompts.
Both the judge (GPT-4.1-nano) and oracle (GPT-5) use identical prompts; the only difference is the underlying model. Temperature = 0.0 for judge, 1.0 for oracle (model constraint).
System prompt: “You are an AI evaluator. Rate responses from 0 to 100. Always provide a score, even if the response is incomplete or truncated.”
User prompt template:
<task>
Evaluate the quality of the AI assistant’s response to the user’s question.
Score from 0-100 based on relevance, helpfulness, clarity, and completeness.
If the response is truncated, score based on what is provided.
</task>
<exchange>
<question>{prompt}</question>
<answer>{response}</answer>
</exchange>
<instruction>
Provide your evaluation score (0-100):
</instruction>
Appendix F Efficient Budget Allocation
While CJE provides calibration-aware confidence intervals for any sample configuration, practitioners often face a constrained optimization problem: minimizing estimation error subject to a fixed financial or latency budget. Here we derive the optimal allocation between cheap surrogate scores ($n$) and expensive oracle labels ($m$).
Proposition 5 (Square Root Allocation Law).
Let $c_{S}$ and $c_{Y}$ be the marginal costs of surrogate scoring and oracle labeling, respectively. Assume the variance decomposes as
$V_{\mathrm{total}}(n,m)\;\approx\;\frac{\sigma^{2}_{\mathrm{eval}}}{n}+\frac{\sigma^{2}_{\mathrm{cal}}}{m},\qquad B=c_{S}n+c_{Y}m.$
The allocation $(n^{*},m^{*})$ minimizing $V_{\mathrm{total}}$ subject to budget $B$ satisfies:
$\boxed{\frac{m^{*}}{n^{*}}=\sqrt{\frac{c_{S}}{c_{Y}}}\cdot\sqrt{\frac{\sigma^{2}_{\mathrm{cal}}}{\sigma^{2}_{\mathrm{eval}}}}}$
with the feasibility constraint $m^{*}\leq n^{*}$ (the oracle slice is a subset of scored examples). If the unconstrained optimum exceeds $n$, set $m^{*}=n$.
Note: This first-order approximation assumes additive variance components with $O(1/m)$ calibration variance. While isotonic regression has nonparametric rates for function estimation, the mean functional $\mathbb{E}[f(S)]$ typically achieves $\sqrt{m}$ rates under smoothness. In practice, estimate $\sigma^{2}_{\mathrm{cal}}\approx m\cdot\widehat{\operatorname{Var}}_{\mathrm{cal}}$ and $\sigma^{2}_{\mathrm{eval}}\approx n\cdot\widehat{\operatorname{Var}}_{\mathrm{eval}}$ from the OUA decomposition, then apply the formula.
Closed-form solution.
Solving the constrained optimization yields:
$n^{*}=\frac{B\,\sqrt{\sigma^{2}_{\mathrm{eval}}/c_{S}}}{\sqrt{c_{S}\sigma^{2}_{\mathrm{eval}}}+\sqrt{c_{Y}\sigma^{2}_{\mathrm{cal}}}},\qquad m^{*}=\frac{B\,\sqrt{\sigma^{2}_{\mathrm{cal}}/c_{Y}}}{\sqrt{c_{S}\sigma^{2}_{\mathrm{eval}}}+\sqrt{c_{Y}\sigma^{2}_{\mathrm{cal}}}}.$
Proof.
We form the Lagrangian $\mathcal{L}$ for minimizing variance subject to cost $B$:
$\mathcal{L}(n,m,\lambda)=\frac{\sigma^{2}_{\mathrm{eval}}}{n}+\frac{\sigma^{2}_{\mathrm{cal}}}{m}+\lambda(c_{S}n+c_{Y}m-B)$
Taking partial derivatives with respect to $n$ and $m$ yields the first-order conditions:
$\frac{\partial\mathcal{L}}{\partial n}=-\frac{\sigma^{2}_{\mathrm{eval}}}{n^{2}}+\lambda c_{S}=0\implies\lambda=\frac{\sigma^{2}_{\mathrm{eval}}}{c_{S}n^{2}}$
$\frac{\partial\mathcal{L}}{\partial m}=-\frac{\sigma^{2}_{\mathrm{cal}}}{m^{2}}+\lambda c_{Y}=0\implies\lambda=\frac{\sigma^{2}_{\mathrm{cal}}}{c_{Y}m^{2}}$
Equating the expressions for $\lambda$:
$\frac{\sigma^{2}_{\mathrm{eval}}}{c_{S}n^{2}}=\frac{\sigma^{2}_{\mathrm{cal}}}{c_{Y}m^{2}}\implies\frac{m^{2}}{n^{2}}=\frac{\sigma^{2}_{\mathrm{cal}}}{\sigma^{2}_{\mathrm{eval}}}\cdot\frac{c_{S}}{c_{Y}}$
Taking the square root yields the ratio result. Substituting into the budget constraint and solving gives the closed-form expressions. ∎
The OUA Marginal Utility Diagnostic.
The optimality condition implies that resources are allocated efficiently when the marginal variance reduction per dollar is equal across both channels. We can express this in terms of the observed OUA Share $\omega=\operatorname{Var}_{\mathrm{cal}}/\operatorname{Var}_{\mathrm{total}}$. At the optimum, the Spend-Balance Rule holds:
$\omega\;=\;\frac{c_{Y}m}{c_{S}n+c_{Y}m}\;=\;\frac{\text{Spend}_{\mathrm{oracle}}}{\text{Spend}_{\mathrm{total}}}.$
That is, the variance share from calibration should equal the budget share spent on oracle labels. Deviations diagnose inefficiency:
  • Under-labeled (Invest in $m$): If $\omega>\frac{c_{Y}m}{c_{S}n+c_{Y}m}$, calibration uncertainty dominates relative to its share of the budget. Action: Shift budget to acquire more oracle labels.
  • Over-labeled (Invest in $n$): If $\omega<\frac{c_{Y}m}{c_{S}n+c_{Y}m}$, evaluation sampling noise dominates. Action: Shift budget to evaluate more prompts with the surrogate.
Worked Example (Arena Benchmark).
Using costs from our experiments: the GPT-4.1-nano judge is approximately $16\times$ cheaper than the GPT-5 oracle, giving a cost ratio $c_{S}/c_{Y}\approx 0.064$. At $n=1000$ prompts with $m=50$ oracle labels (5% oracle fraction), the observed OUA share is $\omega\approx 55\%$.
To find the intrinsic variance ratio, note that $\omega=\frac{\sigma^{2}_{\mathrm{cal}}/m}{\sigma^{2}_{\mathrm{cal}}/m+\sigma^{2}_{\mathrm{eval}}/n}$, which implies:
$\frac{\sigma^{2}_{\mathrm{cal}}}{\sigma^{2}_{\mathrm{eval}}}=\frac{\omega\cdot m}{(1-\omega)\cdot n}=\frac{0.55\times 50}{0.45\times 1000}=0.061$
Applying the Square Root Law:
$\frac{m^{*}}{n^{*}}=\sqrt{0.064}\times\sqrt{0.061}\approx 0.25\times 0.25=6.2\%$
The empirical 5% oracle allocation is close to the variance-optimal 6.2%, confirming that the study’s cost-effective design did not substantially sacrifice precision.
Extension: Multi-Policy Amortization.
When calibrating once to evaluate $P$ policies on the same prompt set, surrogate costs scale as $c_{S}\cdot P\cdot n$ while oracle costs remain $c_{Y}\cdot m$. The optimal ratio becomes $m^{*}/n=\sqrt{P\cdot c_{S}\cdot\sigma^{2}_{\mathrm{cal}}/(c_{Y}\cdot\sigma^{2}_{\mathrm{eval}})}$, favoring larger oracle slices as $P$ grows (calibration cost amortizes across policies, making it relatively cheaper per evaluation).